### v1

# eprint Nisho-2-2025

Abstract
A dark matter axion with mass $m_{a}$ induces an oscillating electric field in a cylindrical sample placed under a magnetic field $B$ parallel to the cylinder axis. When the cylinder is made of a highly electrically conductive material, the induced oscillating current primarily dissipates the axion energy at the surface. In contrast, if the cylinder is composed of a material with low conductivity, e.g. $\sigma=10^{-3}\text{eV}$, the axion energy is dissipated mainly inside the bulk of the cylinder. Within the QCD axion model, the dissipated power $P$ is estimated as $P\simeq 2.8\times 10^{-28}\text{W}g_{\gamma}^{2}\left(L/100\text{cm}\right)\left(R/2\text{cm}\right)^{2}\left(m_{a}/10^{-4}\text{eV}\right)\left(B/10\text{T}\right)^{2}\left(10/\epsilon\right)\left(\rho_{a}/0.3\text{GeVcm}^{-3}\right)$, with $\sigma=10^{-3}$eV, radius $R$, length $L$, electric permittivity $\epsilon=10$ of the cylinder and axion energy density $\rho_{a}$, where $g_{\gamma}(\text{KSVZ})=-0.96$ and $g_{\gamma}(\text{DFSZ})=0.37$. Using an LC circuit appropriately tuned to a quality factor $Q=10^{6}$, the signal-to-noise ratio at low temperature 100100mK and observation time 6060sec. is so large $\sim 10$ that the detection of dark matter axions is feasible in the mass range $m_{a}=10^{-4}\text{-}10^{-3}\text{eV}$.
A central issue in particle physics is to identify phenomena beyond the Standard Model. The axion, proposed as the Nambu–Goldstone boson of Peccei–Quinn symmetry axion ; axion1 ; axion2 , provides a natural solution to the strong CP problem and is also a well-motivated dark matter candidate. It is called as QCD axion. The viable mass window for the QCD axion is tightly constrained to $m_{a}=10^{-6}\mbox{-}10^{-3}\text{eV}$ Wil ; Wil1 ; Wil2 .
Numerous experiments are underway to search for dark matter axions review , most exploiting axion–photon conversion in a strong magnetic field. The induced electromagnetic radiation is expected to be detected with resonant cavitiesadmx ; haystac , superconducting qubitsmoroi , Quantum Hall effect iwazaki ; aiwazaki , e.t.c.
In this letter we propose a new method for axion detection using a cylindrical sample with low electric conductivity, $\sigma=10^{-3}\text{-}10^{-2}\text{eV}$ at low temperature 100100mK. Our target is the mass range $m_{a}=10^{-4}\mbox{-}10^{-3}$eV, which is difficult to detect with resonant cavity.
Dark matter axion generates an oscillating electric field in the presence of a strong magnetic field, which in turn induces an oscillating current in the cylinder. By applying the magnetic field parallel to the cylinder axis, the induced current flows parallel to the external field.
In general, such currents are confined within the skin depth $\delta$. For highly conductive materials, e.g. $\sigma=10^{4}\text{eV}$, the skin depth is extremely small ( $\delta\sim 10^{-5}\text{cm}$ for $m_{a}=10^{-4}$eV ), restricting the current to a thin surface layer. As a result, the effective volume occupied by current flow is small, and the dissipated power from Joule heating is very low to be difficult to detect.
On the other hand, for a cylinder with low conductivity the induced current flows the bulk, and the resulting dissipated power $P$ becomes large enough to be detectable at low temperatures ( $T\sim 100\text{mK}$ ). We show that
$P\simeq 2.8\times 10^{-27}\mbox{W}\Big(\frac{L}{100\mbox{cm}}\Big)\Big(\frac{R}{2\mbox{cm}}\Big)^{2}\Big(\frac{B_{0}}{10\mbox{T}}\Big)^{2}\Big(\frac{m_{a}}{10^{-3}\text{eV}}\Big)\left(\frac{10}{\epsilon}\right)\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)$ (1)
with $\sigma=10m_{a}=10^{-2}$eV, where length L=100L=100cm and radius R=2R=2cm of the cylinder with electric permittivity $\epsilon=10$. The power $P$ depends slightly on $\sigma$ so that even with $\sigma=5\times 10^{-3}$eV, the power is still large $P\simeq 2.2\times 10^{-27}W$. Uniquely, the axion–electromagnetic coupling drives an oscillating current that permeates the entire cylindrical sample, in contrast to ordinary electromagnetic induction where currents are confined to the surface at high frequencies.
By adopting an appropriate conductivity, e.g. $\sigma\sim 5\times 10^{-3}\text{eV}$ at low temperature $\sim 100$mK and radius ( length ) of the cylinder 22cm ( 100100cm ), we show that using LC circuit tuned with a quality factor $Q=10^{6}$, dark matter axion can be probed in the mass range $m_{a}=10^{-4}\text{-}10^{-3}\text{eV}$ with large signal-noise ratio $\sim 10$.
Our method using cylinder with low electric conductivity is also effective for so called, dark photon. The electromagnetic coupling with the dark photon gives rise to similar effect to the one in the axion that oscillating electric current is induced in the whole of cylinder.
The main difficulty in detecting axion dark matter lies in its extremely weak coupling to electromagnetic radiation or ordinary matter ( electrons or nucleons ). In particular, the interaction between the axion field $a(t,\vec{x})$ and the electromagnetic field is described by
$L_{a\gamma\gamma}=g_{a\gamma\gamma}a(t,\vec{x})\vec{E}\cdot\vec{B},$ (2)
where $\vec{E}$ and $\vec{B}$ denote the electric and magnetic fields, respectively. The coupling constant is $g_{a\gamma\gamma}=g_{\gamma}\alpha/\pi f_{a}$, with the fine structure constant $\alpha\simeq 1/137$, the axion decay constant $f_{a}$, and the relation $m_{a}f_{a}\simeq 6\times 10^{-6}\text{eV}\times 10^{12}\text{GeV}$ for the QCD axion model. The model dependent coefficient is $g_{\gamma}\simeq 0.37$ for the DFSZ model dfsz ; dfsz1 and $g_{\gamma}\simeq-0.96$ for the KSVZ model ksvz ; ksvz1 .
For a classical axion field representing dark matter, the interaction term $g_{a\gamma\gamma}a(t,\vec{x})$ is extremely small. Assuming that dark matter consists entirely of axions, the local energy density of the dark matter axion is
$\rho_{d}=m_{a}^{2}\overline{a(t,\vec{x})^{2}}=\frac{1}{2}m_{a}^{2}a_{0}^{2}\simeq 0.3\text{GeV}/\text{cm}^{3},$ (3)
where the overline denotes time averaging. This yields an effective CP-violating interaction of order $g_{a\gamma\gamma}a(t,\vec{x})\sim 10^{-21}$, essentially independent of the QCD axion mass. Consequently, the axion-induced electric field in vacuum under an external magnetic field $B_{0}$ is extremely weak, of order $\sim g_{a\gamma\gamma}aB_{0}$.
In this letter we show that, for a cylindrical sample in Fig.1 with low electrical conductivity, the Joule heating induced by axion dark matter can be large enough to be detectable at low temperatures $T\sim 100\text{mK}$. In this case, the dissipation occurs throughout the bulk of the cylinder, while for a good conductor with large conductivity, the dissipation is confined to a thin surface layer.
We focus on the axion mass range $m_{a}=10^{-4}\text{-}10^{-3}\text{eV}$, using a cylinder of length $L=100\text{cm}$, radius $R\geq 10/m_{a}$, and conductivity $\sigma=10m_{a}=10^{-3}\text{-}10^{-2}\text{eV}$. In particular, we fix the radius R=2R=2cm thought out the paper as a practical value. As we show soon later, the condition of $\sigma m_{a}=\epsilon$ maximize the power dissipated inside the cylinder with electric permittivity $\epsilon$. Such a matter with low conductivity may be realized at low temperatures $\sim 100\text{mK}$, for instance by using semiconductors with suitable impurity doping.
A strong external magnetic field $\vec{B}_{0}$ is applied parallel to the cylinder axis, so that the system is axially symmetric. $\vec{B}_{0}=(0,0,B_{0})$ in cylindrical coordinates $(\rho,\theta,z)$, with $\rho=0$ at the center and $\rho=R$ at the surface of the cylinder. In the presence of dark matter axion, the magnetic field induces an oscillating current parallel to $\vec{B}_{0}$. This current produces microwave radiation with frequencies corresponding to axion masses in the range $m_{a}=10^{-4}\text{-}10^{-3}\text{eV}$.
We calculate the axion-induced electric field to obtain oscillation electric current and its dissipation power $P$.
Figure 1: cylinder sample with length $L$ and radius $R$ under external magnetic field $\vec{B}_{0}$
In order to obtain oscillating electric current $\vec{J}=\sigma\vec{E}^{\prime}$, we solve the Maxwell equations involving axion effect,
$\vec{\partial}\cdot(\epsilon\vec{E}^{\prime}+g_{a\gamma\gamma}a(t,\vec{x})\vec{B})$ $=0$ $,\quad\vec{\partial}\times\Big(\vec{B}-g_{a\gamma\gamma}a(t,\vec{x})\vec{E}^{\prime}\Big)-\partial_{t}\Big(\epsilon\vec{E}^{\prime}+g_{a\gamma\gamma}a(t,\vec{x})\vec{B}\Big)=\vec{J},$ (4)
$\vec{\partial}\cdot\vec{B}$ $=0$ $,\quad\vec{\partial}\times\vec{E}^{\prime}+\partial_{t}\vec{B}=0$ (5)
where electric permittivity $\epsilon$ and electric current $\vec{J}$ are non vanishing inside the cylinder, while $\epsilon=1$ and $\vec{J}=0$ in vacuum. We have assumed trivial permeability $\mu=1$ of the cylinder.
The electric field $\vec{E}^{\prime}$ induced by the axion effect is much small, being of the order of $g_{a\gamma\gamma}aB_{0}$, while $\vec{B}=\vec{B}_{0}+\vec{B}^{\prime}$ with external magnetic field $\vec{B}_{0}=(0,0,B_{0})$. $\vec{B}^{\prime}$ is induced by the axion and associated with $\vec{E}^{\prime}$.
It is easy to obtain the following equation of electric field $\vec{E}^{\prime}$ using cylindrical coordinate.
$\big(\partial_{\rho}^{2}+\frac{1}{\rho}\partial_{\rho}+\epsilon m_{a}^{2}+i\sigma m_{a}\big)\vec{E}^{\prime}=-m_{a}^{2}g_{a\gamma\gamma}a(t)\vec{B}_{0}$ (6)
assuming $a(t)\propto\exp(-im_{a}t)$. The solution $\vec{E}^{\prime}=(0,0,E^{\prime})$ is
$E^{\prime}=d(t)J_{0}(bm_{a}\rho)+\frac{m_{a}^{2}E_{a}}{\epsilon m_{a}^{2}+im_{a}\sigma}$ (7)
with $E_{a}\equiv-g_{a\gamma\gamma}a(t)B_{0}$, where $d(t)\propto\exp(-im_{a}T)$ is a constant determined by boundary conditions at $\rho=R$. The corresponding magnetic field $\vec{B}^{\prime}=(0,B^{\prime},0)$ is given by solving $\partial_{t}B^{\prime}=\partial_{\rho}E^{\prime}$,
$B^{\prime}=-ib\,d(t)J_{1}(bm_{a}\rho)$ (8)
In the above expression, $J_{0,1}(x)$ denotes Bessel function of the first kind and is chosen because of its finiteness of $E^{\prime}(\rho)$ at $\rho=0$. The constant $b$ is given by
$b\equiv(\epsilon^{2}+y^{2})^{1/4}\exp(i\theta/2)\quad\mbox{with}\quad\theta=\cos^{-1}\Big(\frac{\epsilon}{\sqrt{\epsilon^{2}+y^{2}}}\Big)$ (9)
with $y\equiv\sigma/m_{a}$.
These solutions $E^{\prime}$ and $B^{\prime}$ represent electric and magnetic fields inside the cylinder with radius $R$. Similarly, we can find solutions $E_{v}$ and $B_{v}$ outside the cylinder with the conditions $\vec{J}=0$ and $\epsilon=1$ in the Maxwell equations.
$E_{v}=\tilde{d}(t)H_{0}^{(1)}(m_{a}\rho)+E_{a}\quad\mbox{and}\quad B_{v}=-i\tilde{d}(t)H_{1}^{(1)}(m_{a}\rho)$ (10)
with $H_{0,1}^{(1)}$ Hankel functions of the first kind, where $\tilde{d}\propto\exp(-im_{a}t)$ is a constant determined by boundary conditions at $\rho=R$. The Hankel function of the first kind is chosen because the radiations described by $E_{v}$ and $B_{v}$ are outgoing waves; $E_{v}(B_{v})\sim\exp(-im_{a}t+im_{a}\rho)$ as $\rho\to\infty$.
In order to determine the constants $d(t)$ and $\tilde{d}(t)$, we impose boundary conditions at the surface $\rho=R$ of the cylinder such that $\epsilon E^{\prime}(\rho=R)=E_{v}(\rho=R)$ and $B^{\prime}(\rho=R)=B_{v}(\rho=R)$. Then, we have
$\epsilon d(t)J_{0}(bm_{a}R)+\frac{\epsilon m_{a}^{2}E_{a}}{\epsilon m_{a}^{2}+im_{a}\sigma}=\tilde{d}(t)H_{0}^{(1)}(m_{a}R)+E_{a}\quad\mbox{and}\quad b\,d(t)J_{1}(bm_{a}R)=\tilde{d}(t)H_{1}^{(1)}(m_{a}R)$ (11)
Therefore, by solving the equations for $d(t)$ and $\tilde{d}(t)$, we have
$d(t)=\frac{iy}{\epsilon+iy}\frac{E_{a}H_{1}^{(1)}(x)}{\epsilon J_{0}(bx)H_{1}^{(1)}(x)-bJ_{1}(bx)H_{0}^{(1)}(x)},\quad\tilde{d}(t)=d(t)\Bigg(\frac{bJ_{1}(bx)}{H_{1}^{(1)}(x)}\Bigg),$ (12)
with $x=m_{a}R$.
Thus, the oscillating electric and magnetic fields $E^{\prime}$ and $B^{\prime}$ inside the cylinder are
$E^{\prime}(\rho)$ $=$ $E_{a}\Big(\frac{1}{\epsilon+iy}+\frac{iy}{\epsilon+iy}\frac{H_{1}^{(1)}(m_{a}R)J_{0}(bm_{a}\rho)}{\epsilon J_{0}(bm_{a}R)H_{1}^{(1)}(m_{a}R)-bJ_{1}(bm_{a}R)H_{0}^{(1)}(m_{a}R)}\Big)$ (13)
$B^{\prime}(\rho)$ $=$ $E_{a}\Big(\frac{by}{\epsilon+iy}\frac{E_{a}H_{1}^{(1)}(m_{a}R)J_{1}(bm_{a}\rho)}{\epsilon J_{0}(bm_{a}R)H_{1}^{(1)}(m_{a}R)-bJ_{1}(bm_{a}R)H_{0}^{(1)}(m_{a}R)}\Big).$ (14)
They oscillate as $E^{\prime}(B^{\prime})\propto\exp(-im_{a}t)$. It is easy to see that the first term of the electric field, $E^{\prime}(\rho)$, exists throughout the bulk of the cylinder, whereas the second term is confined to its surface; $J_{0}(bm_{a}\rho)/J_{0,1}(bm_{a}R)\propto\exp\big((R-\rho)m_{a}\text{Im}(b)\big)$ where $\text{Im}(O)$ denotes imaginary part of $O$.
Uniquely, the axion–electromagnetic coupling drives an oscillating electric field that permeates the entire cylindrical sample, in contrast to ordinary electromagnetic induction where the electric field is confined to the surface at high frequencies. A similar fact is also present in the case of dark photonkishimoto .
As shown later, the second term dominates the power dissipation in the limit of high conductivity $(\sigma\sim 10^{4}\text{eV})$, while the first term becomes dominant for low conductivity $(\sigma\sim 10^{-3}\text{eV})$.
Using the formula of the electric field $E^{\prime}$, we obtain oscillating electric current $J(\rho)=\sigma Re(E^{\prime}(\rho))$ and corresponding its power $P$ averaged over the period $2\pi/m_{a}$,
$P=\int_{0}^{R}\overline{J(\rho)(Re(E^{\prime}(\rho)))}L2\pi\rho d\rho=\int_{0}^{R}\sigma|E^{\prime}(\rho)|^{2}L\pi\rho d\rho$ (15)
where $Re(O)$ denotes real part of the quantity $O$.
When we put $z=\rho/R$, the formula $P$ is rewritten such that
$P=\sigma L\pi\int_{0}^{1}|E^{\prime}(Rz)|^{2}R^{2}zdz=\frac{\pi L|E_{a}|^{2}}{m_{a}}\int_{0}^{1}yx^{2}|U(x,y,z)|^{2}zdz,$ (16)
where $U(x,y,z)$ is
$U(x,y,z)\equiv\frac{1}{\epsilon+iy}+\frac{iy}{\epsilon+iy}\frac{H_{1}^{(1)}(x)J_{0}(bxz)}{\epsilon J_{0}(bx)H_{1}^{(1)}(x)-bJ_{1}(bx)H_{0}^{(1)}(x)}$ (17)
We remind that $y\equiv\sigma/m_{a}$ and $x\equiv m_{a}R$ and $b=b(y)$ is the function of $y$. $P$ is a complicated function in $x$ and $y$, or $m_{a}$ and $R$. But we show below that $P$ becomes a simple function in $x$ and $y$ when we take the limit of large conductivity, $\sigma\sim 10^{4}$eV or small $\sigma\sim 10^{-3}$eV.
It is easy to confirm that in the limit of $\sigma\to\infty$ ( $y\to\infty$ ), $P\propto m_{a}\rho_{a}V(B_{0}/m_{a}M)^{2}|H_{1}^{(1)}(x)/H_{0}^{(1)}(x)|^{2}$ with $M\equiv\pi f_{a}m_{a}/(g_{\gamma}\alpha)$ and the volume $V=2\pi RL\delta$, within which the electric current flows. $\delta=\sqrt{2/m_{a}\sigma}$ denotes skin depth. The current flows only in the surface with depth $\delta$ of the cylinder. It is coincide with the previous result kishimoto . In the limit, the main contribution comes from the second term in $U$. We note that $|H_{1}^{(1)}(x)/H_{0}^{(1)}(x)|^{2}\simeq 1$ for $x>5$.
The skin depth for general $\sigma$ is given by $\big(m_{a}\mbox{Im}(b(y))\big)^{-1}=\Big(m_{a}(\epsilon^{2}+(\sigma/m_{a})^{2})^{1/4}\sin(\theta/2)\Big)^{-1}>\delta=\sqrt{2/m_{a}\sigma}$. In this paper, assuming typical value $\epsilon=10$ of semiconductor, we take low electric conductivity $\sigma\sim\epsilon m_{a}\sim 10m_{a}$ ( $y\sim 10$ ) at temperature $\sim 100$mK and consider the mass region $m_{a}=10^{-3}\mbox{-}10^{-4}\mbox{eV}$. The matter may be fabricated of for instance a semiconductor with appropriate doping of impurities.
In such a matter with low electric conductivity, we find that the main contribution to $P$ comes from the first term in $U$. Actually, the following quantities are monotonically increasing function in $x$,
$\int_{0}^{1}$ $yx^{2}dzz\Bigg|\frac{iy}{\epsilon+iy}\frac{H_{1}^{(1)}(x)J_{0}(bxz)}{\epsilon J_{0}(bx)H_{1}^{(1)}(x)-bJ_{1}(bx)H_{0}^{(1)}(x)}\Bigg|^{2}=0.01\text{-}0.09\,\,\,\mbox{for}\,\,\,x=1\text{-}10$ (18)
$\int_{0}^{1}$ $yx^{2}dzz\Big|\frac{1}{\epsilon+iy}\Big|^{2}=0.025\text{-}2.5\,\,\,\mbox{for}\,\,\,x=1\text{-}10.$ (19)
with $y=10$ and $\epsilon=10$. We find that for each $x$, the quantity in the second equation is larger than the one in the first equation. Especially, it is more than an order of magnitude for $x\geq 10$. Further, smaller $y$ ( <10<10 ) leads to much larger discrepancy between the quantities. Therefore, we may write
$P=\frac{\pi L|E_{a}|^{2}}{m_{a}}\int_{0}^{1}yx^{2}|U(x,y,z)|^{2}zdz\sim\frac{\pi L|E_{a}|^{2}}{2m_{a}}\Big(\frac{yx^{2}}{\epsilon^{2}+y^{2}}\Big).$ (20)
$P$ takes the maximal value at $y=\sigma/m_{a}=\epsilon$ and $P(\epsilon,y=\epsilon)\propto\epsilon^{-1}$. Thus, it is favorable to take the conductivity $\sigma=\epsilon m_{a}$. Because we do not know the value of the axion mass $m_{a}$, the value $\sigma=\epsilon m_{a}$ is unknown. But when we put $\sigma=\epsilon m_{a}$ in the formula, the dependence of $P$ on $m_{a}$ becomes simple; $P\propto m_{a}$. In the actual search of the axion mass in the range $10^{-4}\text{-}10^{-3}$eV, when we use semiconductor with $\epsilon\sim 10$, it is sufficient for the axion detection to take the value $\sigma\sim 5\times 10^{-3}$eV. Although it does not lead to the maximal power $P$, the predicted signal-noise ratios are large enough for the detection. Hereafter we take $\sigma=\epsilon m_{a}$ and $\epsilon=10$ for simplicity.
Numerically, we find
$P(\sigma=10^{-3}\text{eV})\simeq 2.8\times 10^{-28}\mbox{W}g_{\gamma}^{2}\Big(\frac{L}{100\mbox{cm}}\Big)\Big(\frac{R}{2\mbox{cm}}\Big)^{2}\Big(\frac{B_{0}}{10\mbox{T}}\Big)^{2}\Big(\frac{m_{a}}{10^{-4}\mbox{eV}}\Big)\Big(\frac{10}{\epsilon}\Big)\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)$ (21)
with $\epsilon=10$, $y=10$ ( $\sigma=10m_{a}$ ) and $x\simeq 10$ ( R=2R=2cm ). Obviously, $P$ in eq.(20) is proportional to $R^{2}=x^{2}/m_{a}^{2}$, that is, the cross-sectional area $\pi R^{2}$ of the cylinder. The energy dissipation takes place in the bulk volume $\propto R^{2}L$, not merely in the surface, resulting in substantially enhanced dissipation power $P$.
For comparison, we present the power $P$ when the conductivity is much high, $\sigma=10^{4}$eV,
$P(\sigma=10^{4}\text{eV})$ $\simeq$ $|E_{a}|^{2}\frac{\pi m_{a}\delta RL}{2}\Bigg|\frac{H_{1}^{(1)}(x)}{H_{0}^{(1)}(x)}\Bigg|^{2}$ (22)
$\simeq$ $8.0\times 10^{-32}\mbox{W}g_{\gamma}^{2}\Big(\frac{L}{100\mbox{cm}}\Big)\Big(\frac{R}{2\mbox{cm}}\Big)\sqrt{\frac{m_{a}}{10^{-4}\mbox{eV}}}\sqrt{\frac{10^{4}\mbox{eV}}{\sigma}}\Big(\frac{B_{0}}{10\mbox{T}}\Big)^{2}\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)$ (23)
with $y=\sigma/m_{a}=10^{8}$ and $x=m_{a}R\simeq 10$. It is generated in the surface with the skin depth $\delta\sim 10^{-5}$cm of the cylinder. We find that the power $P(\sigma=10^{-3}\text{eV})$ with low conductivity, is three order of magnitude larger than $P_{h}(\sigma=10^{4}\text{eV})$ with high conductivity. That is, $P(\sigma=10^{-3}\text{eV})>10^{3}P(\sigma=10^{4}\text{eV})$. It is the reason why we use such a cylinder with low conductivity $\sigma\sim 10^{-3}$eV.
We compare $P$ with thermal noise $P_{t}=T\delta\omega/2\pi$ with the width $\delta\omega=10^{-6}m_{a}$ of the microwave frequency,
$P_{t}\simeq 3.3\times 10^{-20}W\Big(\frac{T}{100\rm mK}\Big)\Big(\frac{m_{a}}{10^{-4}\rm eV}\Big)$ (24)
We have
$\frac{P(\sigma=10^{-3}\text{eV})}{P_{t}}\simeq 8.5\times 10^{-9}g_{\gamma}^{2}\Big(\frac{T}{100\rm mK}\Big)^{-1}\Big(\frac{L}{100\mbox{cm}}\Big)\Big(\frac{R}{2\mbox{cm}}\Big)^{2}\Big(\frac{B_{0}}{10\mbox{T}}\Big)^{2}\Big(\frac{10}{\epsilon}\Big)\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)$ (25)
We use LC circuit with resistance $R_{c}$, inductance $L_{c}$ and capacitance $C_{c}$ to detect the power $P$ induced by the axion. When we tune the circuit with resonant frequency $\omega=\sqrt{1/L_{c}C_{c}}=m_{a}$ and a quality factor $Q$ such as $Q=R_{c}^{-1}\sqrt{L_{c}/C_{c}}=10^{6}$, the signal to noise ( S/N ) ratio is given by
$\frac{P(\sigma=10^{-3}\text{eV})}{P_{t}}Q\sqrt{\frac{\delta\omega t_{ob}}{2\pi}}\simeq 10\,g_{\gamma}^{2}\Big(\frac{m_{a}}{10^{-4}\rm eV}\Big)^{1/2}\Big(\frac{T}{100\rm mK}\Big)^{-1}\Big(\frac{L}{100\mbox{cm}}\Big)\Big(\frac{R}{2\mbox{cm}}\Big)^{2}\Big(\frac{B_{0}}{10\mbox{T}}\Big)^{2}\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)\sqrt{\frac{t_{ob}}{60\rm s}}$ (26)
with $y=10$ and $x\simeq 10$, where $t_{ob}$ denotes the observation time.
When we change electric conductivity $\sigma=10^{-3}$eV slightly such as $\sigma=5\times 10^{-4}$eV, the S/N ratio $\simeq 8.2$. The dependence on the conductivity is very weak. But, when axion mass is smaller such as $m_{a}=10^{-5}$eV, it is favorable to use smaller $\sigma=10^{-4}$eV to maximize $P$. Then, it leads to
$\frac{P(\sigma=10^{-4}\text{eV})}{P_{t}}Q\sqrt{\frac{\delta\omega t_{ob}}{2\pi}}\simeq 3.2g_{\gamma}^{2}\,\Big(\frac{m_{a}}{10^{-5}\rm eV}\Big)^{1/2}\Big(\frac{T}{100\rm mK}\Big)^{-1}\Big(\frac{L}{100\mbox{cm}}\Big)\Big(\frac{R}{2\mbox{cm}}\Big)^{2}\Big(\frac{B_{0}}{10\mbox{T}}\Big)^{2}\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)\sqrt{\frac{t_{ob}}{60\rm s}},$ (27)
with $y=10$ and $x\simeq 1$.
Furthermore, when axion mass is larger such as $m_{a}=10^{-3}$eV, it is favorable to use $\sigma=10^{-2}$eV to maximize $P$. It leads to
$\frac{P(\sigma=10^{-2}\text{eV})}{P_{t}}Q\sqrt{\frac{\delta\omega t_{ob}}{2\pi}}\simeq 32\,g_{\gamma}^{2}\,\Big(\frac{m_{a}}{10^{-3}\rm eV}\Big)^{1/2}\Big(\frac{T}{100\rm mK}\Big)^{-1}\Big(\frac{L}{100\mbox{cm}}\Big)\Big(\frac{R}{2\mbox{cm}}\Big)^{2}\Big(\frac{B_{0}}{10\mbox{T}}\Big)^{2}\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)\sqrt{\frac{t_{ob}}{60\rm s}},$ (28)
with $y=10$ and $x\simeq 100$. Even when $\sigma=5\times 10^{-3}$eV, the S/N ratio is $\simeq 26$. We find that the dependence on the conductivity is also weak.
It turns out that we can search the region of the axion mass $m_{a}=10^{-4}\text{-}10^{-3}$eV with large S/N ratio. In particular, even by fixing $\sigma=5\times 10^{-3}$eV and R=2R=2cm, the S/N ratios are $\simeq 8.2$ for $m_{a}=10^{-4}$eV and $\simeq 26$ for $m_{a}=10^{-3}$eV. The dependence of S/N ratio on the conductivity $\sigma$ is weak so that even if the conductivity of the sample is not uniform in the large sample ( L=1L=1m and R=2R=2cm ), the signal of the axion dark matter could be found in the method.
Obviously, the method of the axion detection using cylinder sample with small electric conductivity is very powerful. This is because the axion–electromagnetic coupling induces an oscillating electric current throughout the entire cylindrical sample, producing sufficiently large dissipation power $P$ for detection.
Summarized, we have proposed a method for axion detection using a cylindrical sample with choosing appropriate low conductivity $(\sigma=10^{-3}\text{-}10^{-2}\text{eV})$ sensitive to microwave signals in the frequency range $m_{a}/2\pi=24\text{-}240\text{GHz}$. For such conductivity, the axion-induced electric current penetrates the bulk, yielding substantial energy dissipation and a high signal-to-noise ratio, demonstrating the feasibility of detecting dark matter axions under realistic experimental conditions. The method is also effective for the search of dark photon.

### v2

@@
 A dark matter axion with mass $m_{a}$ induces an oscillating electric field in a cylindrical sample placed under a magnetic field $B$ parallel to the cylinder axis.
-When the cylinder is made of a highly electrically conductive material, the induced oscillating current primarily dissipates the axion energy at the surface.
-In contrast, if the cylinder is composed of a material with low conductivity, e.g.
-$\sigma=10^{-3}\text{eV}$, the axion energy is dissipated mainly inside the bulk of the cylinder.
-Within the QCD axion model, the dissipated power $P$ is estimated as $P\simeq 2.8\times 10^{-28}\text{W}g_{\gamma}^{2}\left(L/100\text{cm}\right)\left(R/2\text{cm}\right)^{2}\left(m_{a}/10^{-4}\text{eV}\right)\left(B/10\text{T}\right)^{2}\left(10/\epsilon\right)\left(\rho_{a}/0.3\text{GeVcm}^{-3}\right)$, with $\sigma=10^{-3}$eV, radius $R$, length $L$, electric permittivity $\epsilon=10$ of the cylinder and axion energy density $\rho_{a}$, where $g_{\gamma}(\text{KSVZ})=-0.96$ and $g_{\gamma}(\text{DFSZ})=0.37$.
-Using an LC circuit appropriately tuned to a quality factor $Q=10^{6}$, the signal-to-noise ratio at low temperature 100100mK and observation time 6060sec.
-is so large $\sim 10$ that the detection of dark matter axions is feasible in the mass range $m_{a}=10^{-4}\text{-}10^{-3}\text{eV}$.
+When the cylinder is made of a highly electrically conductive material, the induced oscillating current flows only at the surface.
+In contrast, if the cylinder is composed of a material with small conductivity, e.g.
+$\sigma=10^{-3}\text{eV}$, the electric current flows inside the bulk of the cylinder.
+Within the QCD axion model, the current $I$ is estimated as $I(m_{a}=10^{-4}\text{eV})\simeq 1.9\times 10^{-14}\text{A}g_{\gamma}(R/6\text{cm})^{2}(\sigma/10^{-3}\text{eV})(B/10\text{T})(10/\epsilon)(\rho_{a}/0.3\text{GeVcm}^{-3})$ with radius $R$, length $L$, permittivity $\epsilon=10$ of the cylinder and axion energy density $\rho_{a}$, where $g_{\gamma}$ is model dependent parameter; $g_{\gamma}(\text{KSVZ})=-0.96$ and $g_{\gamma}(\text{DFSZ})=0.37$.
+Using an LC circuit with a quality factor $Q=10$, the signal-to-noise ratio at temperature T=100T=100mK and observation time $t_{ob}=60$second is $\sim 0.5(Q/10)\sqrt{(m_{a}/10^{-4}\text{eV})(100\text{mK}/T)(t_{ob}/60\text{s})}$.
+The detection of dark matter axions is feasible in the mass range $m_{a}=10^{-4}\text{-}10^{-3}\text{eV}$.
 A central issue in particle physics is to identify phenomena beyond the Standard Model.
@@
 The induced electromagnetic radiation is expected to be detected with resonant cavitiesadmx ; haystac , superconducting qubitsmoroi , Quantum Hall effect iwazaki ; aiwazaki , e.t.c.
-In this letter we propose a new method for axion detection using a cylindrical sample with low electric conductivity, $\sigma=10^{-3}\text{-}10^{-2}\text{eV}$ at low temperature 100100mK.
+In this letter we propose a new method for axion detection using a cylindrical sample with small electrical conductivity, $\sigma=10^{-3}\text{-}10^{-2}\text{eV}$ at low temperature 100100mK.
 Our target is the mass range $m_{a}=10^{-4}\mbox{-}10^{-3}$eV, which is difficult to detect with resonant cavity.
@@
 For highly conductive materials, e.g.
-$\sigma=10^{4}\text{eV}$, the skin depth is extremely small ( $\delta\sim 10^{-5}\text{cm}$ for $m_{a}=10^{-4}$eV ), restricting the current to a thin surface layer.
-As a result, the effective volume occupied by current flow is small, and the dissipated power from Joule heating is very low to be difficult to detect.
-On the other hand, for a cylinder with low conductivity the induced current flows the bulk, and the resulting dissipated power $P$ becomes large enough to be detectable at low temperatures ( $T\sim 100\text{mK}$ ).
-We show that
-$P\simeq 2.8\times 10^{-27}\mbox{W}\Big(\frac{L}{100\mbox{cm}}\Big)\Big(\frac{R}{2\mbox{cm}}\Big)^{2}\Big(\frac{B_{0}}{10\mbox{T}}\Big)^{2}\Big(\frac{m_{a}}{10^{-3}\text{eV}}\Big)\left(\frac{10}{\epsilon}\right)\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)$ (1)
-with $\sigma=10m_{a}=10^{-2}$eV, where length L=100L=100cm and radius R=2R=2cm of the cylinder with electric permittivity $\epsilon=10$.
-The power $P$ depends slightly on $\sigma$ so that even with $\sigma=5\times 10^{-3}$eV, the power is still large $P\simeq 2.2\times 10^{-27}W$.
-Uniquely, the axion–electromagnetic coupling drives an oscillating current that permeates the entire cylindrical sample, in contrast to ordinary electromagnetic induction where currents are confined to the surface at high frequencies.
+$\sigma=10^{4}\text{eV}$, the skin depth is extremely small ( $\delta\sim 10^{-5}\text{cm}$ for $m_{a}=10^{-4}$eV ), restricting the current to a thin surface layer due to the induction effect.
+Additionally, electric field induced in the conducting material is suppressed by the factor $\sqrt{m_{a}/\sigma}\sim 10^{-4}$, compared with the one induced in vacuum.
+As a result, the electric current is not large enough to be detectable.
+On the other hand, the two suppression factors do not arise for electromagnetic fields interacting with axion, Especially, for the cylinder sample with small conductivity, the induced current mainly flows the bulk, and the resulting electric current $I$ becomes large enough to be detectable at low temperatures ( $T\sim 100\text{mK}$ ).
+We show that when the axion mass $m_{a}=10^{-4}$eV,
+$I\simeq 2.1\times 10^{-15}\text{A}g_{\gamma}\Big(\frac{R}{2\text{cm}}\Big)^{2}\Big(\frac{\sigma}{10^{-3}\text{eV}}\Big)\Big(\frac{B_{0}}{10\text{T}}\Big)\left(\frac{10}{\epsilon}\right)\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)$ (1)
+with $\sigma=10m_{a}=10^{-3}$eV, where $R$ denotes the radius of the cylinder with permittivity $\epsilon=10$.
+The current is maximized by choosing electrical conductivity $\sigma=\epsilon m_{a}$, for instance, $\sigma=10^{-3}\text{eV}(m_{a}/10^{-4}\text{eV})$ when $\epsilon=10$.
+$I$ is proportional to the surface area $R^{2}$.
+The oscillating current flows the entire cylindrical sample, in contrast to ordinary electromagnetic induction where currents with high frequencies are confined to the surface.
+In the actual measurement we compare the current $I$ with current $I_{n}$ of thermal noise.
+When the noise current $I_{n}$ is smaller, the real signal current $I$ can be observed more easily.
+Thus, it is favorable that the noise $I_{n}$ is small.
+We adopt cylinder sample with small conductivity, e.g.
+$\sigma=10^{-3}$eV, which causes large resistance $R_{c}$ of the cylinder.
+Because $I_{n}\propto 1/\sqrt{R_{c}}$, it is expected that the thermal noise $I_{n}$ is small for large $R_{c}$.
+It is an advantage of using the sample with small conductivity $\sigma=10^{-3}$eV for the axion detection.
+As we show later, we have
+$\frac{I}{I_{n}}\simeq 0.43\times 10^{-4}g_{\gamma}\sqrt{\frac{100\rm mK}{T}}\sqrt{\frac{\sigma}{10^{-3}\text{eV}}}\sqrt{\frac{10^{-4}\text{eV}}{m_{a}}}\sqrt{\frac{L}{100\mbox{cm}}}\Big(\frac{R}{2\mbox{cm}}\Big)\Big(\frac{B_{0}}{10\mbox{T}}\Big)\left(\frac{10}{\epsilon}\right)\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)$ (2)
+with length $L$ of the cylinder.
 By adopting an appropriate conductivity, e.g.
-$\sigma\sim 5\times 10^{-3}\text{eV}$ at low temperature $\sim 100$mK and radius ( length ) of the cylinder 22cm ( 100100cm ), we show that using LC circuit tuned with a quality factor $Q=10^{6}$, dark matter axion can be probed in the mass range $m_{a}=10^{-4}\text{-}10^{-3}\text{eV}$ with large signal-noise ratio $\sim 10$.
-Our method using cylinder with low electric conductivity is also effective for so called, dark photon.
-The electromagnetic coupling with the dark photon gives rise to similar effect to the one in the axion that oscillating electric current is induced in the whole of cylinder.
+$\sigma\sim 5\times 10^{-3}\text{eV}$ at low temperature $\sim 20$mK and radius ( length ) of the cylinder 66cm ( 100100cm ), we show that using parallel LC circuit tuned with a quality factor $Q=10$, dark matter axion can be probed in the mass range $m_{a}=10^{-4}\text{-}10^{-3}\text{eV}$ with signal-noise ratio of the electric current bigger than $10$.
+Our method using cylinder with small electrical conductivity is also effective for so called, dark photon.
+The electromagnetic coupling with the dark photon gives rise to similar effect to the one in the QCD axion that oscillating electric current is induced in the whole of cylinder.
 The main difficulty in detecting axion dark matter lies in its extremely weak coupling to electromagnetic radiation or ordinary matter ( electrons or nucleons ).
@@
 Consequently, the axion-induced electric field in vacuum under an external magnetic field $B_{0}$ is extremely weak, of order $\sim g_{a\gamma\gamma}aB_{0}$.
-In this letter we show that, for a cylindrical sample in Fig.1 with low electrical conductivity, the Joule heating induced by axion dark matter can be large enough to be detectable at low temperatures $T\sim 100\text{mK}$.
-In this case, the dissipation occurs throughout the bulk of the cylinder, while for a good conductor with large conductivity, the dissipation is confined to a thin surface layer.
-We focus on the axion mass range $m_{a}=10^{-4}\text{-}10^{-3}\text{eV}$, using a cylinder of length $L=100\text{cm}$, radius $R\geq 10/m_{a}$, and conductivity $\sigma=10m_{a}=10^{-3}\text{-}10^{-2}\text{eV}$.
-In particular, we fix the radius R=2R=2cm thought out the paper as a practical value.
-As we show soon later, the condition of $\sigma m_{a}=\epsilon$ maximize the power dissipated inside the cylinder with electric permittivity $\epsilon$.
-Such a matter with low conductivity may be realized at low temperatures $\sim 100\text{mK}$, for instance by using semiconductors with suitable impurity doping.
+In this letter we show that, for a cylindrical sample in Fig.1 with small electrical conductivity, the electric current induced by axion dark matter can be large enough to be detectable at low temperatures $T\sim 100\text{mK}$.
+In this case, the current flows throughout the bulk of the cylinder, while for a good conductor with large conductivity, the current is confined to a thin surface layer.
+We focus on the axion mass range $m_{a}=10^{-4}\text{-}10^{-3}\text{eV}$, using a cylinder of length $L=100\text{cm}$, radius $R\geq 2$ cm, and conductivity $\sigma=10m_{a}=10^{-3}\text{-}10^{-2}\text{eV}$.
+As we show soon later, the condition of $\sigma/m_{a}=\epsilon$ maximize the electric field induced in the cylinder with permittivity $\epsilon$.
+Such a matter with small conductivity may be realized at low temperatures $\sim 100\text{mK}$, for instance by using semiconductors with impurity doping appropriately.
 A strong external magnetic field $\vec{B}_{0}$ is applied parallel to the cylinder axis, so that the system is axially symmetric.
@@
 This current produces microwave radiation with frequencies corresponding to axion masses in the range $m_{a}=10^{-4}\text{-}10^{-3}\text{eV}$.
-We calculate the axion-induced electric field to obtain oscillation electric current and its dissipation power $P$.
+We calculate the axion-induced electric field $E^{\prime}$ to obtain oscillating electric current.
 Figure 1: cylinder sample with length $L$ and radius $R$ under external magnetic field $\vec{B}_{0}$
@@
 $\vec{\partial}\cdot\vec{B}$ $=0$ $,\quad\vec{\partial}\times\vec{E}^{\prime}+\partial_{t}\vec{B}=0$ (6)
-where electric permittivity $\epsilon$ and electric current $\vec{J}$ are non vanishing inside the cylinder, while $\epsilon=1$ and $\vec{J}=0$ in vacuum.
+where permittivity $\epsilon$ and electric current $\vec{J}$ are non vanishing inside the cylinder, while $\epsilon=1$ and $\vec{J}=0$ in vacuum.
 We have assumed trivial permeability $\mu=1$ of the cylinder.
@@
 They oscillate as $E^{\prime}(B^{\prime})\propto\exp(-im_{a}t)$.
-It is easy to see that the first term of the electric field, $E^{\prime}(\rho)$, exists throughout the bulk of the cylinder, whereas the second term is confined to its surface; $J_{0}(bm_{a}\rho)/J_{0,1}(bm_{a}R)\propto\exp\big((R-\rho)m_{a}\text{Im}(b)\big)$ where $\text{Im}(O)$ denotes imaginary part of $O$.
-Uniquely, the axion–electromagnetic coupling drives an oscillating electric field that permeates the entire cylindrical sample, in contrast to ordinary electromagnetic induction where the electric field is confined to the surface at high frequencies.
+Obviously, the first term of the electric field, $E^{\prime}(\rho)$, exists throughout the bulk of the cylinder, whereas the second term is confined to its surface; $J_{0}(bm_{a}\rho)/J_{0,1}(bm_{a}R)\propto\exp\big((R-\rho)m_{a}\text{Im}(b)\big)$ as $\rho\to\infty$ where $\text{Im}(O)$ denotes imaginary part of $O$.
+Uniquely, the axion–electromagnetic coupling drives an oscillating electric field that permeates the entire cylindrical sample, in contrast to ordinary electromagnetic induction, where the high frequency electric field is confined to the surface.
 A similar fact is also present in the case of dark photonkishimoto .
-As shown later, the second term dominates the power dissipation in the limit of high conductivity $(\sigma\sim 10^{4}\text{eV})$, while the first term becomes dominant for low conductivity $(\sigma\sim 10^{-3}\text{eV})$.
+As shown later, the second term dominates the electric power ( Joule heat ) in the limit of large conductivity $(\sigma\sim 10^{4}\text{eV})$, while the first term becomes dominant for small conductivity $(\sigma\sim 10^{-3}\text{eV})$.
 Using the formula of the electric field $E^{\prime}$, we obtain oscillating electric current $J(\rho)=\sigma Re(E^{\prime}(\rho))$ and corresponding its power $P$ averaged over the period $2\pi/m_{a}$,
@@
 But we show below that $P$ becomes a simple function in $x$ and $y$ when we take the limit of large conductivity, $\sigma\sim 10^{4}$eV or small $\sigma\sim 10^{-3}$eV.
-It is easy to confirm that in the limit of $\sigma\to\infty$ ( $y\to\infty$ ), $P\propto m_{a}\rho_{a}V(B_{0}/m_{a}M)^{2}|H_{1}^{(1)}(x)/H_{0}^{(1)}(x)|^{2}$ with $M\equiv\pi f_{a}m_{a}/(g_{\gamma}\alpha)$ and the volume $V=2\pi RL\delta$, within which the electric current flows.
+We note that electric field is written such that $E^{\prime}(\rho)=E_{a}U$.
+It is easy to confirm that in the limit of $\sigma\to\infty$ ( $y\to\infty$ ), $P\propto m_{a}\rho_{a}V(B_{0}/m_{a}M)^{2}|H_{1}^{(1)}(x)/H_{0}^{(1)}(x)|^{2}$ with $M\equiv\pi f_{a}/(g_{\gamma}\alpha)$ and the volume $V=2\pi RL\delta$, within which the electric current flows.
 $\delta=\sqrt{2/m_{a}\sigma}$ denotes skin depth.
-The current flows only in the surface with depth $\delta$ of the cylinder.
+The current flows only in the surface with the depth $\delta$ of the cylinder.
 It is coincide with the previous result kishimoto .
@@
 We note that $|H_{1}^{(1)}(x)/H_{0}^{(1)}(x)|^{2}\simeq 1$ for $x>5$.
-The skin depth for general $\sigma$ is given by $\big(m_{a}\mbox{Im}(b(y))\big)^{-1}=\Big(m_{a}(\epsilon^{2}+(\sigma/m_{a})^{2})^{1/4}\sin(\theta/2)\Big)^{-1}>\delta=\sqrt{2/m_{a}\sigma}$.
-In this paper, assuming typical value $\epsilon=10$ of semiconductor, we take low electric conductivity $\sigma\sim\epsilon m_{a}\sim 10m_{a}$ ( $y\sim 10$ ) at temperature $\sim 100$mK and consider the mass region $m_{a}=10^{-3}\mbox{-}10^{-4}\mbox{eV}$.
-The matter may be fabricated of for instance a semiconductor with appropriate doping of impurities.
-In such a matter with low electric conductivity, we find that the main contribution to $P$ comes from the first term in $U$.
+The skin depth for general $\sigma$ represented by the second term of $U(x,y,z)$ in eq(18) is given by $\big(m_{a}\mbox{Im}(b(y))\big)^{-1}=\Big(m_{a}(\epsilon^{2}+(\sigma/m_{a})^{2})^{1/4}\sin(\theta/2)\Big)^{-1}>\delta=\sqrt{2/m_{a}\sigma}$.
+In this paper, assuming typical value $\epsilon=10$ in semiconductor, we take small electric conductivity $\sigma\sim\epsilon m_{a}\sim 10m_{a}$ ( $y\sim 10$ ) and consider the mass region $m_{a}=10^{-3}\mbox{-}10^{-4}\mbox{eV}$.
+Then, the skin depth $\big(m_{a}\mbox{Im}(b(y))\big)^{-1}\simeq 0.1\text{cm}(10^{-4}\text{eV}/m_{a})$, smaller than the radius R=2R=2cm examined in our paper.
+On the other hand, the first term of $U$ is present in the whole of the cylinder, not confined at the surface.
+In such a cylinder with small electrical conductivity, we find that the main contribution to $P$ comes from the first term in $U$.
 Actually, the following quantities are monotonically increasing function in $x$,
@@
 But when we put $\sigma=\epsilon m_{a}$ in the formula, the dependence of $P$ on $m_{a}$ becomes simple; $P\propto m_{a}$.
-In the actual search of the axion mass in the range $10^{-4}\text{-}10^{-3}$eV, when we use semiconductor with $\epsilon\sim 10$, it is sufficient for the axion detection to take the value $\sigma\sim 5\times 10^{-3}$eV.
+In the actual search of the axion mass in the range $10^{-4}\text{-}10^{-3}$eV, it is sufficient for the axion detection to take the value $\sigma\sim 5\times 10^{-3}$eV with $\epsilon\sim 10$.
 Although it does not lead to the maximal power $P$, the predicted signal-noise ratios are large enough for the detection.
 Hereafter we take $\sigma=\epsilon m_{a}$ and $\epsilon=10$ for simplicity.
-Numerically, we find
+Numerically, noting $|E_{a}|=|g_{a\gamma\gamma}a(t)B_{0}|\simeq 1.6\times 10^{-36}\text{GeV}^{2}g_{\gamma}(B_{0}/10\text{T})$, we find
 $P(\sigma=10^{-3}\text{eV})\simeq 2.8\times 10^{-28}\mbox{W}g_{\gamma}^{2}\Big(\frac{L}{100\mbox{cm}}\Big)\Big(\frac{R}{2\mbox{cm}}\Big)^{2}\Big(\frac{B_{0}}{10\mbox{T}}\Big)^{2}\Big(\frac{m_{a}}{10^{-4}\mbox{eV}}\Big)\Big(\frac{10}{\epsilon}\Big)\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)$ (22)
 with $\epsilon=10$, $y=10$ ( $\sigma=10m_{a}$ ) and $x\simeq 10$ ( R=2R=2cm ).
-Obviously, $P$ in eq.(20) is proportional to $R^{2}=x^{2}/m_{a}^{2}$, that is, the cross-sectional area $\pi R^{2}$ of the cylinder.
+Obviously, $P$ in eq.(21) is proportional to $R^{2}=x^{2}/m_{a}^{2}$, that is, the cross section $\pi R^{2}$ of the cylinder.
 The energy dissipation takes place in the bulk volume $\propto R^{2}L$, not merely in the surface, resulting in substantially enhanced dissipation power $P$.
@@
 It is generated in the surface with the skin depth $\delta\sim 10^{-5}$cm of the cylinder.
-We find that the power $P(\sigma=10^{-3}\text{eV})$ with low conductivity, is three order of magnitude larger than $P_{h}(\sigma=10^{4}\text{eV})$ with high conductivity.
+We find that the power $P(\sigma=10^{-3}\text{eV})$ with small conductivity, is three order of magnitude larger than $P(\sigma=10^{4}\text{eV})$ with large conductivity.
 That is, $P(\sigma=10^{-3}\text{eV})>10^{3}P(\sigma=10^{4}\text{eV})$.
-It is the reason why we use such a cylinder with low conductivity $\sigma\sim 10^{-3}$eV.
-We compare $P$ with thermal noise $P_{t}=T\delta\omega/2\pi$ with the width $\delta\omega=10^{-6}m_{a}$ of the microwave frequency,
-$P_{t}\simeq 3.3\times 10^{-20}W\Big(\frac{T}{100\rm mK}\Big)\Big(\frac{m_{a}}{10^{-4}\rm eV}\Big)$ (24)
-We have
-$\frac{P(\sigma=10^{-3}\text{eV})}{P_{t}}\simeq 8.5\times 10^{-9}g_{\gamma}^{2}\Big(\frac{T}{100\rm mK}\Big)^{-1}\Big(\frac{L}{100\mbox{cm}}\Big)\Big(\frac{R}{2\mbox{cm}}\Big)^{2}\Big(\frac{B_{0}}{10\mbox{T}}\Big)^{2}\Big(\frac{10}{\epsilon}\Big)\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)$ (25)
-We use LC circuit with resistance $R_{c}$, inductance $L_{c}$ and capacitance $C_{c}$ to detect the power $P$ induced by the axion.
-When we tune the circuit with resonant frequency $\omega=\sqrt{1/L_{c}C_{c}}=m_{a}$ and a quality factor $Q$ such as $Q=R_{c}^{-1}\sqrt{L_{c}/C_{c}}=10^{6}$, the signal to noise ( S/N ) ratio is given by
-$\frac{P(\sigma=10^{-3}\text{eV})}{P_{t}}Q\sqrt{\frac{\delta\omega t_{ob}}{2\pi}}\simeq 10\,g_{\gamma}^{2}\Big(\frac{m_{a}}{10^{-4}\rm eV}\Big)^{1/2}\Big(\frac{T}{100\rm mK}\Big)^{-1}\Big(\frac{L}{100\mbox{cm}}\Big)\Big(\frac{R}{2\mbox{cm}}\Big)^{2}\Big(\frac{B_{0}}{10\mbox{T}}\Big)^{2}\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)\sqrt{\frac{t_{ob}}{60\rm s}}$ (26)
-with $y=10$ and $x\simeq 10$, where $t_{ob}$ denotes the observation time.
-When we change electric conductivity $\sigma=10^{-3}$eV slightly such as $\sigma=5\times 10^{-4}$eV, the S/N ratio $\simeq 8.2$.
-The dependence on the conductivity is very weak.
-But, when axion mass is smaller such as $m_{a}=10^{-5}$eV, it is favorable to use smaller $\sigma=10^{-4}$eV to maximize $P$.
-Then, it leads to
-$\frac{P(\sigma=10^{-4}\text{eV})}{P_{t}}Q\sqrt{\frac{\delta\omega t_{ob}}{2\pi}}\simeq 3.2g_{\gamma}^{2}\,\Big(\frac{m_{a}}{10^{-5}\rm eV}\Big)^{1/2}\Big(\frac{T}{100\rm mK}\Big)^{-1}\Big(\frac{L}{100\mbox{cm}}\Big)\Big(\frac{R}{2\mbox{cm}}\Big)^{2}\Big(\frac{B_{0}}{10\mbox{T}}\Big)^{2}\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)\sqrt{\frac{t_{ob}}{60\rm s}},$ (27)
-with $y=10$ and $x\simeq 1$.
-Furthermore, when axion mass is larger such as $m_{a}=10^{-3}$eV, it is favorable to use $\sigma=10^{-2}$eV to maximize $P$. It leads to
-$\frac{P(\sigma=10^{-2}\text{eV})}{P_{t}}Q\sqrt{\frac{\delta\omega t_{ob}}{2\pi}}\simeq 32\,g_{\gamma}^{2}\,\Big(\frac{m_{a}}{10^{-3}\rm eV}\Big)^{1/2}\Big(\frac{T}{100\rm mK}\Big)^{-1}\Big(\frac{L}{100\mbox{cm}}\Big)\Big(\frac{R}{2\mbox{cm}}\Big)^{2}\Big(\frac{B_{0}}{10\mbox{T}}\Big)^{2}\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)\sqrt{\frac{t_{ob}}{60\rm s}},$ (28)
-with $y=10$ and $x\simeq 100$. Even when $\sigma=5\times 10^{-3}$eV, the S/N ratio is $\simeq 26$. We find that the dependence on the conductivity is also weak.
-It turns out that we can search the region of the axion mass $m_{a}=10^{-4}\text{-}10^{-3}$eV with large S/N ratio.
-In particular, even by fixing $\sigma=5\times 10^{-3}$eV and R=2R=2cm, the S/N ratios are $\simeq 8.2$ for $m_{a}=10^{-4}$eV and $\simeq 26$ for $m_{a}=10^{-3}$eV.
-The dependence of S/N ratio on the conductivity $\sigma$ is weak so that even if the conductivity of the sample is not uniform in the large sample ( L=1L=1m and R=2R=2cm ), the signal of the axion dark matter could be found in the method.
-Obviously, the method of the axion detection using cylinder sample with small electric conductivity is very powerful.
-This is because the axion–electromagnetic coupling induces an oscillating electric current throughout the entire cylindrical sample, producing sufficiently large dissipation power $P$ for detection.
-Summarized, we have proposed a method for axion detection using a cylindrical sample with choosing appropriate low conductivity $(\sigma=10^{-3}\text{-}10^{-2}\text{eV})$ sensitive to microwave signals in the frequency range $m_{a}/2\pi=24\text{-}240\text{GHz}$.
-For such conductivity, the axion-induced electric current penetrates the bulk, yielding substantial energy dissipation and a high signal-to-noise ratio, demonstrating the feasibility of detecting dark matter axions under realistic experimental conditions.
-The method is also effective for the search of dark photon.
+It is the reason why we use such a cylinder with small conductivity $\sigma\sim 10^{-3}$eV.
+There are two reasons of the small power $P(\sigma=10^{4}\text{eV})$ in the cylinder.
+One is the suppression of electric field inside conductor with large conductivity $\sigma=10^{4}$eV, that is, $\sqrt{m_{a}/\sigma}|E_{a}|$ aiwazaki ; kishimoto compared with $|E_{a}|/2\epsilon$ in material with small conductivity $\sigma=10^{-3}$eV.
+The other one is the presence of the skin depth $\delta=\sqrt{2/m_{a}\sigma}$.
+The area $2\pi R\delta$ of the current flowing is much smaller than that $\pi R^{2}$ in material with low conductivity.
+Therefore, even if large conductivity $\sigma$ amplifies electric current by the factor $\sigma$, both the suppression of electric field by the factor $\sqrt{m_{a}/\sigma}$ and the area of current flow by the factor $\sqrt{1/m_{a}\sigma}$, cause the power $P(\sigma=10^{4}\text{eV})$ smaller than the power $P(\sigma=10^{-3}\text{eV})$.
+It is roughly,
+$\frac{P(\sigma_{h}=10^{4}\text{eV})}{P(\sigma=10^{-3}\text{eV})}\sim\frac{R\delta\sigma_{h}(\sqrt{m_{a}/\sigma_{h}})^{2}}{R^{2}\sigma(1/\epsilon^{2})}=\frac{\delta m_{a}\epsilon^{2}}{R\sigma}\sim 10^{-3}$ (25)
+with R=1R=1cm, $\epsilon=10$ and $m_{a}=10^{-4}$eV. It is coincident with the above calculations in eq(22) and eq(23).
+The large power $P(\sigma=10^{-3}\text{eV})$ arises mainly from large electric current $I$ flowing in the bulk of the cylinder. The electric current $I$ is,
+$I(\sigma=10^{-3}\text{eV})=\frac{\sigma|E_{a}|S}{2\epsilon}\simeq 2.1\times 10^{-15}\text{A}g_{\gamma}\Big(\frac{R}{2\text{cm}}\Big)^{2}\Big(\frac{\sigma}{10^{-3}\text{eV}}\Big)\Big(\frac{B_{0}}{10\text{T}}\Big),$ (26)
+where the electric field $E=Re(E^{\prime})=|E_{a}|/2\epsilon=|E_{a}|/20$ is the one in the first term with $y=10$ in eq(14). We detect the axion by measuring the large electric current $I$.
+The point causing the large current $I$ is that electric field $E$ inside of the cylinder with small conductivity, e.g.
+$\sigma=10^{-3}$eV has no suppression factor $\sqrt{m_{a}/\sigma}\simeq 10^{-4}$, which is present when the cylinder with large conductivity $\sigma=10^{4}$eV.
+Additionally, the current flows in the whole of the cylinder, not just in the surface.
+The facts make the large electric current $I$ to be observable.
+There is another point which makes axion detection feasible.
+In the actual observation, we need to take into account thermal noise ( Johnson-Nyquist noise ).
+The electric current $I_{n}$ by the thermal noise is given by $I_{n}=\sqrt{2T\delta\omega/\pi R_{c}}$ with temperature T=100T=100mK and the width $\delta\omega$ of the frequency; $\delta\omega=10^{-6}m_{a}=10^{-10}(m_{a}/10^{-4}\text{eV})$eV.
+$R_{c}=L/(\sigma S)$ with $S=\pi R^{2}$ denotes electric resistance of the cylinder; $R_{c}=0.16(L/100\text{cm})(2\text{cm}/R)^{2}$.
+Then we find
+$I_{n}=4.9\times 10^{-11}\text{A}\sqrt{\Big(\frac{T}{100\rm mK}\Big)\Big(\frac{m_{a}}{10^{-4}\rm eV}\Big)\Big(\frac{\sigma}{10^{-3}\text{eV}}\Big)\Big(\frac{100\text{cm}}{L}\Big)\Big(\frac{R}{2\text{cm}}\Big)^{2}}.$ (27)
+Because the noise $I_{n}\propto 1/\sqrt{R_{c}}$, the noise is small for the large resistance $R_{c}\simeq 0.16$ of the cylinder.
+It should be stressed that the small noise $I_{n}$ is caused by small electrical conductivity $\sigma=10^{-3}$eV of the cylinder.
+Thus, we obtain the ratio of the current $I$ to the noise $I_{n}$,
+$\frac{I(\sigma=10^{-3}\text{eV})}{I_{n}(\sigma=10^{-3}\text{eV})}\simeq 0.43\times 10^{-4}g_{\gamma}\sqrt{\frac{100\rm mK}{T}}\sqrt{\frac{\sigma}{10^{-3}\text{eV}}}\sqrt{\frac{10^{-4}\text{eV}}{m_{a}}}\sqrt{\frac{L}{100\mbox{cm}}}\Big(\frac{R}{2\mbox{cm}}\Big)\Big(\frac{B_{0}}{10\mbox{T}}\Big)\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)$ (28)
+In order to amplify the above ratio $I/I_{n}$, we use, for instance, parallel LC circuit with resistance $R_{c}$, inductance $L_{c}$ and capacitance $C_{c}$.
+( The resistance $R_{c}$ in the parallel LC circuit is almost equal to the one of the cylinder.
+) When we tune the circuit with resonant frequency $\omega=\sqrt{1/L_{c}C_{c}}=m_{a}$ and a quality factor $Q$ such as $Q=R_{c}\sqrt{C_{c}/L_{c}}$, the above ratio is amplified such as $Q(I/I_{n})$.
+In actual experiment, it is difficult to have large quality factor $Q$ such as $Q>10^{3}$.
+So we assume $Q=10$ as a realistic value and measure the electric current in the coil of the LC circuit.
+Furthermore, in the measurement of the oscillating current, we take the frequency width $\delta\omega=10^{-6}m_{a}$ and observation time $60$ seconds.
+Then, the actual noise is reduced as $I_{n}/\sqrt{\delta\omega\delta t_{ob}/(2\pi)}\simeq I_{n}/(1.2\times 10^{3}\sqrt{\frac{m_{a}}{10^{-4}\text{eV}}}\sqrt{\frac{t_{ob}}{60s}})$.
+It follows that the signal-noise ( SN ) ratio is
+$\frac{I(\sigma=10^{-3}\text{eV})}{I_{n}(\sigma=10^{-3}\text{eV})}Q\sqrt{\frac{\delta\omega t_{ob}}{2\pi}}\simeq 0.51\,g_{\gamma}\Big(\frac{Q}{10}\Big)\sqrt{\frac{100\rm mK}{T}}\sqrt{\frac{t_{ob}}{60\rm\,s}}\sqrt{\frac{\sigma}{10^{-3}\text{eV}}}\sqrt{\frac{L}{100\mbox{cm}}}\Big(\frac{R}{2\mbox{cm}}\Big)\Big(\frac{B_{0}}{10\mbox{T}}\Big)\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)$ (29)
+with $y=10$ and $x\simeq 10$, where $t_{ob}=60$ s denotes the observation time.
+Although the SN ratio does not explicitly involve the axion mass $m_{a}$, it implicitly depends on $m_{a}$ because $\sigma$ is taken as $\sigma=10m_{a}$.
+When we searches $m_{a}=10^{-3}$eV, it is favorable to have $\sigma=10^{-2}$eV.
+Then, the SN ratio is $\simeq 1.6$.
+Therefore, it turns out that by the observation of the electric current $I$ with appropriate choices of, for instance, temperature T=20T=20mK and the radius R=6R=6cm, and observation time $t_{ob}=600$ s, SN ratio is larger than $10$.
+Without such a choice of the parameters, we can obtain large SN ratio as far as we tune quality factor $Q$ bigger than $10$.
+We point out that when we search much larger axion mass $m_{a}>10^{-3}$eV, it is easier to detect such axions This is because electric current $I\propto\sigma$ and SN ratio $\propto\sqrt{\sigma}$ are much larger than those of $m_{a}=10^{-3}$eV with the appropriate choice of conductivity $\sigma=10m_{a}$.
+We find that the method of the axion detection using cylinder sample with small electrical conductivity is very powerful.
+This is because the axion–electromagnetic coupling induces an oscillating electric field $|E_{a}|/2\epsilon$ throughout the entire cylindrical sample, producing sufficiently large electric current $I\propto\sigma(|E_{a}|/\epsilon)^{2}R^{2}$ for the detection.
+Furthermore, the sample with small electrical conductivity has small thermal noise compared with the one with large electrical conductivity.
+Summarized, we have proposed a method for axion detection with a cylindrical sample choosing appropriate small conductivity $(\sigma=10^{-3}\text{-}10^{-2}\text{eV})$ sensitive to microwave signals in the frequency range $m_{a}/2\pi=24\text{-}240\text{GHz}$.
+With such conductivity, the axion induces large electric current flowing the bulk of the cylinder and causes a high signal-to-noise ratio of the current.
+Therefore, the method demonstrates the feasibility of detecting dark matter axions under realistic experimental conditions.
+It is also effective for the search of dark photon.

### v3

@@
 Abstract
-A dark matter axion with mass $m_{a}$ induces an oscillating electric field in a cylindrical sample placed under a magnetic field $B$ parallel to the cylinder axis.
+A dark matter axion with mass $m_{a}$ induces an oscillating electric field in a cylindrical sample placed under a magnetic field $B_{0}$ parallel to the cylinder axis.
 When the cylinder is made of a highly electrically conductive material, the induced oscillating current flows only at the surface.
@@
 $\sigma=10^{-3}\text{eV}$, the electric current flows inside the bulk of the cylinder.
-Within the QCD axion model, the current $I$ is estimated as $I(m_{a}=10^{-4}\text{eV})\simeq 1.9\times 10^{-14}\text{A}g_{\gamma}(R/6\text{cm})^{2}(\sigma/10^{-3}\text{eV})(B/10\text{T})(10/\epsilon)(\rho_{a}/0.3\text{GeVcm}^{-3})$ with radius $R$, length $L$, permittivity $\epsilon=10$ of the cylinder and axion energy density $\rho_{a}$, where $g_{\gamma}$ is model dependent parameter; $g_{\gamma}(\text{KSVZ})=-0.96$ and $g_{\gamma}(\text{DFSZ})=0.37$.
-Using an LC circuit with a quality factor $Q=10$, the signal-to-noise ratio at temperature T=100T=100mK and observation time $t_{ob}=60$second is $\sim 0.5(Q/10)\sqrt{(m_{a}/10^{-4}\text{eV})(100\text{mK}/T)(t_{ob}/60\text{s})}$.
-The detection of dark matter axions is feasible in the mass range $m_{a}=10^{-4}\text{-}10^{-3}\text{eV}$.
+Within the QCD axion model, the power $P$ is estimated as $P\simeq 5.7\times 10^{-27}\mbox{W}g_{\gamma}^{2}\Big(\frac{L}{100\mbox{cm}}\Big)\Big(\frac{R}{6\mbox{cm}}\Big)^{2}\Big(\frac{B_{0}}{15\mbox{T}}\Big)^{2}\Big(\frac{m_{a}}{10^{-4}\mbox{eV}}\Big)\Big(\frac{10}{\epsilon}\Big)\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)$ with radius $R$, length $L$, permittivity $\epsilon=10$ of the cylinder and axion energy density $\rho_{a}$, where $g_{\gamma}$ is model dependent parameter; $g_{\gamma}(\text{KSVZ})=-0.96$ and $g_{\gamma}(\text{DFSZ})=0.37$.
+The detection of dark matter axions in our method may be feasible in the mass range $m_{a}=10^{-4}\text{-}10^{-3}\text{eV}$.
+I introduction
 A central issue in particle physics is to identify phenomena beyond the Standard Model.
@@
 The induced electromagnetic radiation is expected to be detected with resonant cavitiesadmx ; haystac , superconducting qubitsmoroi , Quantum Hall effect iwazaki ; aiwazaki , e.t.c.
-In this letter we propose a new method for axion detection using a cylindrical sample with small electrical conductivity, $\sigma=10^{-3}\text{-}10^{-2}\text{eV}$ at low temperature 100100mK.
+In this letter we propose a new method for axion detection using a cylindrical sample with small electrical conductivity, $\sigma=10^{-3}\text{-}10^{-2}\text{eV}$ at low temperature 2020mK.
 Our target is the mass range $m_{a}=10^{-4}\mbox{-}10^{-3}$eV, which is difficult to detect with resonant cavity.
@@
 As a result, the electric current is not large enough to be detectable.
-On the other hand, the two suppression factors do not arise for electromagnetic fields interacting with axion, Especially, for the cylinder sample with small conductivity, the induced current mainly flows the bulk, and the resulting electric current $I$ becomes large enough to be detectable at low temperatures ( $T\sim 100\text{mK}$ ).
+On the other hand, the two suppression factors do not arise in materials with small conductivity such as semiconductor, Especially, for the cylinder sample with small conductivity $\sigma=10^{-3}$eV, the induced current mainly flows the bulk, and the resulting electric current $I$ and power $P$ become large.
+It would be detectable at low temperatures ( $T\sim 10\text{mK}$ ).
 We show that when the axion mass $m_{a}=10^{-4}$eV,
-$I\simeq 2.1\times 10^{-15}\text{A}g_{\gamma}\Big(\frac{R}{2\text{cm}}\Big)^{2}\Big(\frac{\sigma}{10^{-3}\text{eV}}\Big)\Big(\frac{B_{0}}{10\text{T}}\Big)\left(\frac{10}{\epsilon}\right)\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)$ (1)
-with $\sigma=10m_{a}=10^{-3}$eV, where $R$ denotes the radius of the cylinder with permittivity $\epsilon=10$.
-The current is maximized by choosing electrical conductivity $\sigma=\epsilon m_{a}$, for instance, $\sigma=10^{-3}\text{eV}(m_{a}/10^{-4}\text{eV})$ when $\epsilon=10$.
-$I$ is proportional to the surface area $R^{2}$.
-The oscillating current flows the entire cylindrical sample, in contrast to ordinary electromagnetic induction where currents with high frequencies are confined to the surface.
-In the actual measurement we compare the current $I$ with current $I_{n}$ of thermal noise.
-When the noise current $I_{n}$ is smaller, the real signal current $I$ can be observed more easily.
-Thus, it is favorable that the noise $I_{n}$ is small.
-We adopt cylinder sample with small conductivity, e.g.
-$\sigma=10^{-3}$eV, which causes large resistance $R_{c}$ of the cylinder.
-Because $I_{n}\propto 1/\sqrt{R_{c}}$, it is expected that the thermal noise $I_{n}$ is small for large $R_{c}$.
-It is an advantage of using the sample with small conductivity $\sigma=10^{-3}$eV for the axion detection.
-As we show later, we have
-$\frac{I}{I_{n}}\simeq 0.43\times 10^{-4}g_{\gamma}\sqrt{\frac{100\rm mK}{T}}\sqrt{\frac{\sigma}{10^{-3}\text{eV}}}\sqrt{\frac{10^{-4}\text{eV}}{m_{a}}}\sqrt{\frac{L}{100\mbox{cm}}}\Big(\frac{R}{2\mbox{cm}}\Big)\Big(\frac{B_{0}}{10\mbox{T}}\Big)\left(\frac{10}{\epsilon}\right)\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)$ (2)
-with length $L$ of the cylinder.
-By adopting an appropriate conductivity, e.g.
-$\sigma\sim 5\times 10^{-3}\text{eV}$ at low temperature $\sim 20$mK and radius ( length ) of the cylinder 66cm ( 100100cm ), we show that using parallel LC circuit tuned with a quality factor $Q=10$, dark matter axion can be probed in the mass range $m_{a}=10^{-4}\text{-}10^{-3}\text{eV}$ with signal-noise ratio of the electric current bigger than $10$.
-Our method using cylinder with small electrical conductivity is also effective for so called, dark photon.
-The electromagnetic coupling with the dark photon gives rise to similar effect to the one in the QCD axion that oscillating electric current is induced in the whole of cylinder.
+$I$ $\simeq$ $2.8\times 10^{-14}\text{A}g_{\gamma}\Big(\frac{R}{6\text{cm}}\Big)^{2}\Big(\frac{\sigma}{10^{-3}\text{eV}}\Big)\Big(\frac{B_{0}}{15\text{T}}\Big)\left(\frac{10}{\epsilon}\right)\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)$ (1)
+$P(\sigma=10^{-3}\text{eV})$ $\simeq$ $5.7\times 10^{-27}\mbox{W}g_{\gamma}^{2}\Big(\frac{L}{100\mbox{cm}}\Big)\Big(\frac{R}{6\mbox{cm}}\Big)^{2}\Big(\frac{B_{0}}{15\mbox{T}}\Big)^{2}\Big(\frac{m_{a}}{10^{-4}\mbox{eV}}\Big)\Big(\frac{10}{\epsilon}\Big)\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)$ (2)
+with the choice $\sigma=10m_{a}=10^{-3}$eV, where $R(L)$ denotes the radius ( length ) of the cylinder with permittivity $\epsilon=10$.
+The power is maximized by choosing electrical conductivity such as $\sigma=\epsilon m_{a}$, for instance, $\sigma=10^{-3}\text{eV}(m_{a}/10^{-4}\text{eV})$ when $\epsilon=10$.
+We note that $I$ and $P(\sigma=10^{-3}\text{eV})$ are proportional to the surface area $R^{2}$.
+The oscillating current flows the entire cylindrical sample, in contrast to ordinal metal with large conductivity, in which electric currents with high frequencies are confined to the surface due to electromagnetic induction.
+Indeed, when the conductivity is large, for instance, $\sigma=10^{4}$eV, the power is given by
+$P(\sigma=10^{4}\text{eV})\simeq 5.4\times 10^{-31}\mbox{W}g_{\gamma}^{2}\Big(\frac{L}{100\mbox{cm}}\Big)\Big(\frac{R}{6\mbox{cm}}\Big)\sqrt{\frac{m_{a}}{10^{-4}\mbox{eV}}}\sqrt{\frac{10^{4}\mbox{eV}}{\sigma}}\Big(\frac{B}{15\mbox{T}}\Big)^{2}\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)$ (3)
+Obviously, the power $P(\sigma=10^{4}\text{eV})$ proportional to $R$ is fourth order of magnitude smaller than $P(\sigma=10^{-3}\text{eV})$.
+In the observation, the power $P(\sigma=10^{-3}\text{eV})$ should be compared with thermal noise $P_{n}=T\delta\omega/2\pi$ with $\delta\omega=10^{-6}m_{a}$,
+$\frac{P(\sigma=10^{-3}\text{eV})}{P_{n}}\simeq 8.6\times 10^{-7}g_{\gamma}^{2}\Big(\frac{20\text{mK}}{T}\Big)\Big(\frac{L}{100\mbox{cm}}\Big)\Big(\frac{R}{6\text{cm}}\Big)^{2}\Big(\frac{B}{15\text{T}}\Big)^{2}\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)$ (4)
+The noise is further reduced by considering a frequency width $\delta\omega=10^{-6}m_{a}$ and an observation time $\delta t_{ob}=100$seconds.
+In this case, the noise decreases by the factor $\sqrt{\delta\omega\delta t_{ob}/2\pi}\simeq 1.6\times 10^{3}$, yielding $P(\sigma=10^{-3}\text{eV})\sqrt{\delta\omega\delta t_{ob}/2\pi}/P_{n}\simeq 1.3\times 10^{-3}$.
+Although the details of power detection are beyond the scope of this paper, we expect the ratio to be sufficiently large for the power ( $P(\sigma=10^{-3},\text{eV})$ ) to become detectable in the near future.
+Our method, which employs a cylinder with low electrical conductivity, is also applicable to the detection of the so-called dark photon.
+The electromagnetic coupling of the dark photon produces an effect similar to that of the axion.
+II axion
 The main difficulty in detecting axion dark matter lies in its extremely weak coupling to electromagnetic radiation or ordinary matter ( electrons or nucleons ).
@@
 where $\vec{E}$ and $\vec{B}$ denote the electric and magnetic fields, respectively.
-The coupling constant is $g_{a\gamma\gamma}=g_{\gamma}\alpha/\pi f_{a}$, with the fine structure constant $\alpha\simeq 1/137$, the axion decay constant $f_{a}$, and the relation $m_{a}f_{a}\simeq 6\times 10^{-6}\text{eV}\times 10^{12}\text{GeV}$ for the QCD axion model.
+The coupling constant is $g_{a\gamma\gamma}=g_{\gamma}\alpha/\pi f_{a}$, with the fine structure constant $\alpha\simeq 1/137$, the axion decay constant $f_{a}$, and the relation $m_{a}f_{a}\simeq 6\times 10^{-3}\text{GeV}^{2}$ for the QCD axion model.
 The model dependent coefficient is $g_{\gamma}\simeq 0.37$ for the DFSZ model dfsz ; dfsz1 and $g_{\gamma}\simeq-0.96$ for the KSVZ model ksvz ; ksvz1 .
@@
 Consequently, the axion-induced electric field in vacuum under an external magnetic field $B_{0}$ is extremely weak, of order $\sim g_{a\gamma\gamma}aB_{0}$.
-In this letter we show that, for a cylindrical sample in Fig.1 with small electrical conductivity, the electric current induced by axion dark matter can be large enough to be detectable at low temperatures $T\sim 100\text{mK}$.
-In this case, the current flows throughout the bulk of the cylinder, while for a good conductor with large conductivity, the current is confined to a thin surface layer.
-We focus on the axion mass range $m_{a}=10^{-4}\text{-}10^{-3}\text{eV}$, using a cylinder of length $L=100\text{cm}$, radius $R\geq 2$ cm, and conductivity $\sigma=10m_{a}=10^{-3}\text{-}10^{-2}\text{eV}$.
-As we show soon later, the condition of $\sigma/m_{a}=\epsilon$ maximize the electric field induced in the cylinder with permittivity $\epsilon$.
-Such a matter with small conductivity may be realized at low temperatures $\sim 100\text{mK}$, for instance by using semiconductors with impurity doping appropriately.
+III electric power generated in cylinder with small conductivity
+In this letter we show that, for a cylindrical sample in Fig.1 with small electrical conductivity, the electric power induced by axion dark matter can be large and may be detectable at low temperatures $T\sim 10\text{mK}$.
+The current flows throughout the bulk of the cylinder, while for a good conductor with large conductivity, the current is confined to a thin surface layer.
+We focus on the axion mass range $m_{a}=10^{-4}\text{-}10^{-3}\text{eV}$, using a cylinder of length $L=100\text{cm}$, radius $R=6$ cm, and conductivity $\sigma=10m_{a}=10^{-3}\text{-}10^{-2}\text{eV}$.
+As we show soon later, the condition of $\sigma=\epsilon m_{a}$ maximize the electric power induced in the cylinder with permittivity $\epsilon$.
+Such a matter with small conductivity must be realized at low temperatures $\sim 10\text{mK}$, for instance by using semiconductors with impurity doping appropriately.
 A strong external magnetic field $\vec{B}_{0}$ is applied parallel to the cylinder axis, so that the system is axially symmetric.
@@
 $P$ is a complicated function in $x$ and $y$, or $m_{a}$ and $R$.
-But we show below that $P$ becomes a simple function in $x$ and $y$ when we take the limit of large conductivity, $\sigma\sim 10^{4}$eV or small $\sigma\sim 10^{-3}$eV.
+But we show below that $P$ becomes a simple function in $x$ and $y$ when we take the limit of large conductivity, $\sigma\sim 10^{4}$eV or small one $\sigma\sim 10^{-3}$eV.
 We note that electric field is written such that $E^{\prime}(\rho)=E_{a}U$.
-It is easy to confirm that in the limit of $\sigma\to\infty$ ( $y\to\infty$ ), $P\propto m_{a}\rho_{a}V(B_{0}/m_{a}M)^{2}|H_{1}^{(1)}(x)/H_{0}^{(1)}(x)|^{2}$ with $M\equiv\pi f_{a}/(g_{\gamma}\alpha)$ and the volume $V=2\pi RL\delta$, within which the electric current flows.
+It is easy to confirm that in the limit of $\sigma\to\infty$ ( $y\to\infty$ ), $P\propto m_{a}\rho_{a}V(B_{0}/m_{a}M)^{2}|H_{1}^{(1)}(x)/H_{0}^{(1)}(x)|^{2}$ with $M\equiv\pi f_{a}/(g_{\gamma}\alpha)=1/g_{a\gamma\gamma}$ and the volume $V=2\pi RL\delta$, within which the electric current flows.
 $\delta=\sqrt{2/m_{a}\sigma}$ denotes skin depth.
 The current flows only in the surface with the depth $\delta$ of the cylinder.
-It is coincide with the previous result kishimoto .
+It is coincident with the previous result kishimoto .
 In the limit, the main contribution comes from the second term in $U$.
@@
 In this paper, assuming typical value $\epsilon=10$ in semiconductor, we take small electric conductivity $\sigma\sim\epsilon m_{a}\sim 10m_{a}$ ( $y\sim 10$ ) and consider the mass region $m_{a}=10^{-3}\mbox{-}10^{-4}\mbox{eV}$.
-Then, the skin depth $\big(m_{a}\mbox{Im}(b(y))\big)^{-1}\simeq 0.1\text{cm}(10^{-4}\text{eV}/m_{a})$, smaller than the radius R=2R=2cm examined in our paper.
+Then, the skin depth $\big(m_{a}\mbox{Im}(b(y))\big)^{-1}\simeq 0.1\text{cm}(10^{-4}\text{eV}/m_{a})$, smaller than the radius R=6R=6cm examined in our paper.
 On the other hand, the first term of $U$ is present in the whole of the cylinder, not confined at the surface.
@@
 Hereafter we take $\sigma=\epsilon m_{a}$ and $\epsilon=10$ for simplicity.
-Numerically, noting $|E_{a}|=|g_{a\gamma\gamma}a(t)B_{0}|\simeq 1.6\times 10^{-36}\text{GeV}^{2}g_{\gamma}(B_{0}/10\text{T})$, we find
-$P(\sigma=10^{-3}\text{eV})\simeq 2.8\times 10^{-28}\mbox{W}g_{\gamma}^{2}\Big(\frac{L}{100\mbox{cm}}\Big)\Big(\frac{R}{2\mbox{cm}}\Big)^{2}\Big(\frac{B_{0}}{10\mbox{T}}\Big)^{2}\Big(\frac{m_{a}}{10^{-4}\mbox{eV}}\Big)\Big(\frac{10}{\epsilon}\Big)\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)$ (22)
-with $\epsilon=10$, $y=10$ ( $\sigma=10m_{a}$ ) and $x\simeq 10$ ( R=2R=2cm ).
-Obviously, $P$ in eq.(21) is proportional to $R^{2}=x^{2}/m_{a}^{2}$, that is, the cross section $\pi R^{2}$ of the cylinder.
+Numerically, noting $|E_{a}|=|g_{a\gamma\gamma}a(t)B_{0}|\simeq 2.4\times 10^{-36}\text{GeV}^{2}g_{\gamma}(B_{0}/15\text{T})$, we find
+$P(\sigma=10^{-3}\text{eV})\simeq 5.7\times 10^{-27}\mbox{W}g_{\gamma}^{2}\Big(\frac{L}{100\mbox{cm}}\Big)\Big(\frac{R}{6\mbox{cm}}\Big)^{2}\Big(\frac{B_{0}}{15\mbox{T}}\Big)^{2}\Big(\frac{m_{a}}{10^{-4}\mbox{eV}}\Big)\Big(\frac{10}{\epsilon}\Big)\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)$ (24)
+with $\epsilon=10$, $y=10$ ( $\sigma=10m_{a}$ ) and $x\simeq 30$ ( R=6R=6cm ).
+Obviously, $P(\sigma=10^{-3}\text{eV})$ in eq.(24) is proportional to $R^{2}=x^{2}/m_{a}^{2}$, that is, the cross section $\pi R^{2}$ of the cylinder.
 The energy dissipation takes place in the bulk volume $\propto R^{2}L$, not merely in the surface, resulting in substantially enhanced dissipation power $P$.
@@
 $P(\sigma=10^{4}\text{eV})$ $\simeq$ $|E_{a}|^{2}\frac{\pi m_{a}\delta RL}{2}\Bigg|\frac{H_{1}^{(1)}(x)}{H_{0}^{(1)}(x)}\Bigg|^{2}$ (25)
-$\simeq$ $8.0\times 10^{-32}\mbox{W}g_{\gamma}^{2}\Big(\frac{L}{100\mbox{cm}}\Big)\Big(\frac{R}{2\mbox{cm}}\Big)\sqrt{\frac{m_{a}}{10^{-4}\mbox{eV}}}\sqrt{\frac{10^{4}\mbox{eV}}{\sigma}}\Big(\frac{B_{0}}{10\mbox{T}}\Big)^{2}\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)$ (24)
+$\simeq$ $5.4\times 10^{-31}\mbox{W}g_{\gamma}^{2}\Big(\frac{L}{100\mbox{cm}}\Big)\Big(\frac{R}{6\mbox{cm}}\Big)\sqrt{\frac{m_{a}}{10^{-4}\mbox{eV}}}\sqrt{\frac{10^{4}\mbox{eV}}{\sigma}}\Big(\frac{B_{0}}{15\mbox{T}}\Big)^{2}\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)$
 with $y=\sigma/m_{a}=10^{8}$ and $x=m_{a}R\simeq 30$.
 It is generated in the surface with the skin depth $\delta\sim 10^{-5}$cm of the cylinder.
-We find that the power $P(\sigma=10^{-3}\text{eV})$ with small conductivity, is three order of magnitude larger than $P(\sigma=10^{4}\text{eV})$ with large conductivity.
-That is, $P(\sigma=10^{-3}\text{eV})>10^{3}P(\sigma=10^{4}\text{eV})$.
+We find that the power $P(\sigma=10^{-3}\text{eV})$ is fourth order of magnitude larger than $P(\sigma=10^{4}\text{eV})$.
+That is, $P(\sigma=10^{-3}\text{eV})>10^{4}P(\sigma=10^{4}\text{eV})$.
 It is the reason why we use such a cylinder with small conductivity $\sigma\sim 10^{-3}$eV.
@@
 The large power $P(\sigma=10^{-3}\text{eV})$ arises mainly from large electric current $I$ flowing in the bulk of the cylinder. The electric current $I$ is,
-$I(\sigma=10^{-3}\text{eV})=\frac{\sigma|E_{a}|S}{2\epsilon}\simeq 2.1\times 10^{-15}\text{A}g_{\gamma}\Big(\frac{R}{2\text{cm}}\Big)^{2}\Big(\frac{\sigma}{10^{-3}\text{eV}}\Big)\Big(\frac{B_{0}}{10\text{T}}\Big),$ (26)
-where the electric field $E=Re(E^{\prime})=|E_{a}|/2\epsilon=|E_{a}|/20$ is the one in the first term with $y=10$ in eq(14). We detect the axion by measuring the large electric current $I$.
-The point causing the large current $I$ is that electric field $E$ inside of the cylinder with small conductivity, e.g.
+$I(\sigma=10^{-3}\text{eV})=\frac{\sigma|E_{a}|S}{2\epsilon}\simeq 2.8\times 10^{-14}\text{A}g_{\gamma}\Big(\frac{R}{6\text{cm}}\Big)^{2}\Big(\frac{\sigma}{10^{-3}\text{eV}}\Big)\Big(\frac{B_{0}}{15\text{T}}\Big),$ (27)
+where the electric field $E=Re(E^{\prime})=|E_{a}|/2\epsilon=|E_{a}|/20$ is the one in the first term with $y=10$ in eq(16).
+The point causing the large current $I(\sigma=10^{-3}\text{eV})$ is that electric field $E$ inside of the cylinder with small conductivity, e.g.
 $\sigma=10^{-3}$eV has no suppression factor $\sqrt{m_{a}/\sigma}\simeq 10^{-4}$, which is present when the cylinder with large conductivity $\sigma=10^{4}$eV.
 Additionally, the current flows in the whole of the cylinder, not just in the surface.
-The facts make the large electric current $I$ to be observable.
-There is another point which makes axion detection feasible.
-In the actual observation, we need to take into account thermal noise ( Johnson-Nyquist noise ).
-The electric current $I_{n}$ by the thermal noise is given by $I_{n}=\sqrt{2T\delta\omega/\pi R_{c}}$ with temperature T=100T=100mK and the width $\delta\omega$ of the frequency; $\delta\omega=10^{-6}m_{a}=10^{-10}(m_{a}/10^{-4}\text{eV})$eV.
-$R_{c}=L/(\sigma S)$ with $S=\pi R^{2}$ denotes electric resistance of the cylinder; $R_{c}=0.16(L/100\text{cm})(2\text{cm}/R)^{2}$.
-Then we find
-$I_{n}=4.9\times 10^{-11}\text{A}\sqrt{\Big(\frac{T}{100\rm mK}\Big)\Big(\frac{m_{a}}{10^{-4}\rm eV}\Big)\Big(\frac{\sigma}{10^{-3}\text{eV}}\Big)\Big(\frac{100\text{cm}}{L}\Big)\Big(\frac{R}{2\text{cm}}\Big)^{2}}.$ (27)
-Because the noise $I_{n}\propto 1/\sqrt{R_{c}}$, the noise is small for the large resistance $R_{c}\simeq 0.16$ of the cylinder.
-It should be stressed that the small noise $I_{n}$ is caused by small electrical conductivity $\sigma=10^{-3}$eV of the cylinder.
-Thus, we obtain the ratio of the current $I$ to the noise $I_{n}$,
-$\frac{I(\sigma=10^{-3}\text{eV})}{I_{n}(\sigma=10^{-3}\text{eV})}\simeq 0.43\times 10^{-4}g_{\gamma}\sqrt{\frac{100\rm mK}{T}}\sqrt{\frac{\sigma}{10^{-3}\text{eV}}}\sqrt{\frac{10^{-4}\text{eV}}{m_{a}}}\sqrt{\frac{L}{100\mbox{cm}}}\Big(\frac{R}{2\mbox{cm}}\Big)\Big(\frac{B_{0}}{10\mbox{T}}\Big)\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)$ (28)
-In order to amplify the above ratio $I/I_{n}$, we use, for instance, parallel LC circuit with resistance $R_{c}$, inductance $L_{c}$ and capacitance $C_{c}$.
-( The resistance $R_{c}$ in the parallel LC circuit is almost equal to the one of the cylinder.
-) When we tune the circuit with resonant frequency $\omega=\sqrt{1/L_{c}C_{c}}=m_{a}$ and a quality factor $Q$ such as $Q=R_{c}\sqrt{C_{c}/L_{c}}$, the above ratio is amplified such as $Q(I/I_{n})$.
-In actual experiment, it is difficult to have large quality factor $Q$ such as $Q>10^{3}$.
-So we assume $Q=10$ as a realistic value and measure the electric current in the coil of the LC circuit.
-Furthermore, in the measurement of the oscillating current, we take the frequency width $\delta\omega=10^{-6}m_{a}$ and observation time $60$ seconds.
-Then, the actual noise is reduced as $I_{n}/\sqrt{\delta\omega\delta t_{ob}/(2\pi)}\simeq I_{n}/(1.2\times 10^{3}\sqrt{\frac{m_{a}}{10^{-4}\text{eV}}}\sqrt{\frac{t_{ob}}{60s}})$.
-It follows that the signal-noise ( SN ) ratio is
-$\frac{I(\sigma=10^{-3}\text{eV})}{I_{n}(\sigma=10^{-3}\text{eV})}Q\sqrt{\frac{\delta\omega t_{ob}}{2\pi}}\simeq 0.51\,g_{\gamma}\Big(\frac{Q}{10}\Big)\sqrt{\frac{100\rm mK}{T}}\sqrt{\frac{t_{ob}}{60\rm\,s}}\sqrt{\frac{\sigma}{10^{-3}\text{eV}}}\sqrt{\frac{L}{100\mbox{cm}}}\Big(\frac{R}{2\mbox{cm}}\Big)\Big(\frac{B_{0}}{10\mbox{T}}\Big)\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)$ (29)
-with $y=10$ and $x\simeq 10$, where $t_{ob}=60$ s denotes the observation time.
-Although the SN ratio does not explicitly involve the axion mass $m_{a}$, it implicitly depends on $m_{a}$ because $\sigma$ is taken as $\sigma=10m_{a}$.
-When we searches $m_{a}=10^{-3}$eV, it is favorable to have $\sigma=10^{-2}$eV.
-Then, the SN ratio is $\simeq 1.6$.
-Therefore, it turns out that by the observation of the electric current $I$ with appropriate choices of, for instance, temperature T=20T=20mK and the radius R=6R=6cm, and observation time $t_{ob}=600$ s, SN ratio is larger than $10$.
-Without such a choice of the parameters, we can obtain large SN ratio as far as we tune quality factor $Q$ bigger than $10$.
-We point out that when we search much larger axion mass $m_{a}>10^{-3}$eV, it is easier to detect such axions This is because electric current $I\propto\sigma$ and SN ratio $\propto\sqrt{\sigma}$ are much larger than those of $m_{a}=10^{-3}$eV with the appropriate choice of conductivity $\sigma=10m_{a}$.
-We find that the method of the axion detection using cylinder sample with small electrical conductivity is very powerful.
-This is because the axion–electromagnetic coupling induces an oscillating electric field $|E_{a}|/2\epsilon$ throughout the entire cylindrical sample, producing sufficiently large electric current $I\propto\sigma(|E_{a}|/\epsilon)^{2}R^{2}$ for the detection.
-Furthermore, the sample with small electrical conductivity has small thermal noise compared with the one with large electrical conductivity.
+The facts make the large electric current $I(\sigma=10^{-3}\text{eV})\sim 10^{-14}$A.
+In order to observe the power $P(\sigma=10^{-3}\text{eV})$, we need to take into account thermal noise ( Johnson-Nyquist noise ).
+The noise power $P_{n}$ is given such that $P_{n}=T\delta\omega/2\pi$ with $\delta\omega=10^{-6}m_{a}$.
+$\delta\omega$ denotes frequency width in the observation.
+Then we have
+$P_{n}=\frac{T\delta\omega}{2\pi}\simeq 6.6\times 10^{-21}\text{W}\Big(\frac{T}{20\text{mK}}\Big)\Big(\frac{m_{a}}{10^{-4}\text{eV}}\Big).$ (28)
+The noise is reduced by the factor $\sqrt{\delta\omega\delta t_{ob}/2\pi}$ with observation time $\delta t_{ob}$. It yields
+$P_{n}\to\frac{P_{n}}{\sqrt{\frac{\delta\omega\delta t_{ob}}{2\pi}}}\simeq 4.2\times 10^{-24}\text{W}\Big(\frac{T}{20\text{mK}}\Big)\Big(\frac{m_{a}}{10^{-4}\text{eV}}\Big)^{1/2}\Big(\frac{100\,\text{s}}{\delta t_{ob}}\Big)^{-1/2}.$ (29)
+Therefore, the ratio of $P(\sigma=10^{-3}\text{eV})$ to the noise $P_{n}/\sqrt{\delta\omega\delta t_{ob}/2\pi}$ is
+$\frac{P(\sigma=10^{-3}\text{eV})}{P_{n}}\sqrt{\frac{\delta\omega\delta t_{ob}}{2\pi}}\simeq 1.4\times 10^{-3}g_{\gamma}^{2}\Big(\frac{20\text{mK}}{T}\Big)\Big(\frac{L}{100\mbox{cm}}\Big)\Big(\frac{R}{6\mbox{cm}}\Big)^{2}\times$
+$\Big(\frac{B_{0}}{15\mbox{T}}\Big)^{2}\Big(\frac{m_{a}}{10^{-4}\mbox{eV}}\Big)^{1/2}\Big(\frac{10}{\epsilon}\Big)\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)\Big(\frac{\delta t_{ob}}{100\,\text{s}}\Big)^{1/2}.$ (30)
+It is the ratio when $m_{a}=10^{-4}$eV and $\sigma=10^{-3}$eV. Further, when $m_{a}=10^{-3}$eV, taking $\sigma=10m_{a}=10^{-2}$eV, we have
+$\frac{P(\sigma=10^{-2}\text{eV})}{P_{n}}\sqrt{\frac{\delta\omega\delta t_{ob}}{2\pi}}\simeq 4.4\times 10^{-3}g_{\gamma}^{2}\Big(\frac{20\text{mK}}{T}\Big)\Big(\frac{L}{100\mbox{cm}}\Big)\Big(\frac{R}{6\mbox{cm}}\Big)^{2}\times$
+$\Big(\frac{B_{0}}{15\mbox{T}}\Big)^{2}\Big(\frac{m_{a}}{10^{-3}\mbox{eV}}\Big)^{1/2}\Big(\frac{10}{\epsilon}\Big)\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)\Big(\frac{\delta t_{ob}}{100\,\text{s}}\Big)^{1/2}.$ (31)
+To achieve a signal-to-noise ratio greater than 1, the power $P$ must be amplified by approximately three orders of magnitude.
+The method of amplification and its technical implementation are beyond the scope of this paper.
+However, once such amplification is realized, the dark matter axion can be detected using our proposed approach.
+IV conclusion
 Summarized, we have proposed a method for axion detection with a cylindrical sample choosing appropriate small conductivity $(\sigma=10^{-3}\text{-}10^{-2}\text{eV})$ sensitive to microwave signals in the frequency range $m_{a}/2\pi=24\text{-}240\text{GHz}$.
-With such conductivity, the axion induces large electric current flowing the bulk of the cylinder and causes a high signal-to-noise ratio of the current.
+With such conductivity, the axion induces large electric power generated in the bulk of the cylinder and causes a high signal-to-noise ratio.
 Therefore, the method demonstrates the feasibility of detecting dark matter axions under realistic experimental conditions.

### v4

@@
 Within the QCD axion model, the power $P$ is estimated as $P\simeq 5.7\times 10^{-27}\mbox{W}g_{\gamma}^{2}\Big(\frac{L}{100\mbox{cm}}\Big)\Big(\frac{R}{6\mbox{cm}}\Big)^{2}\Big(\frac{B_{0}}{15\mbox{T}}\Big)^{2}\Big(\frac{m_{a}}{10^{-4}\mbox{eV}}\Big)\Big(\frac{10}{\epsilon}\Big)\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)$ with radius $R$, length $L$, permittivity $\epsilon=10$ of the cylinder and axion energy density $\rho_{a}$, where $g_{\gamma}$ is model dependent parameter; $g_{\gamma}(\text{KSVZ})=-0.96$ and $g_{\gamma}(\text{DFSZ})=0.37$.
+The signal of the axion can be detected by observing electric current in parallel LC circuit with quality factor $Q\geq 2$.
 The detection of dark matter axions in our method may be feasible in the mass range $m_{a}=10^{-4}\text{-}10^{-3}\text{eV}$.
@@
 We show that when the axion mass $m_{a}=10^{-4}$eV,
-$I$ $\simeq$ $2.8\times 10^{-14}\text{A}g_{\gamma}\Big(\frac{R}{6\text{cm}}\Big)^{2}\Big(\frac{\sigma}{10^{-3}\text{eV}}\Big)\Big(\frac{B_{0}}{15\text{T}}\Big)\left(\frac{10}{\epsilon}\right)\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)$ (1)
+$I(\sigma=10^{-3}\text{eV})$ $\simeq$ $2.8\times 10^{-14}\text{A}g_{\gamma}\Big(\frac{R}{6\text{cm}}\Big)^{2}\Big(\frac{\sigma}{10^{-3}\text{eV}}\Big)\Big(\frac{B_{0}}{15\text{T}}\Big)\left(\frac{10}{\epsilon}\right)\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)^{1/2}$ (1)
 $P(\sigma=10^{-3}\text{eV})$ $\simeq$ $5.7\times 10^{-27}\mbox{W}g_{\gamma}^{2}\Big(\frac{L}{100\mbox{cm}}\Big)\Big(\frac{R}{6\mbox{cm}}\Big)^{2}\Big(\frac{B_{0}}{15\mbox{T}}\Big)^{2}\Big(\frac{m_{a}}{10^{-4}\mbox{eV}}\Big)\Big(\frac{10}{\epsilon}\Big)\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)$ (2)
@@
 Obviously, the power $P(\sigma=10^{4}\text{eV})$ proportional to $R$ is fourth order of magnitude smaller than $P(\sigma=10^{-3}\text{eV})$.
-In the observation, the power $P(\sigma=10^{-3}\text{eV})$ should be compared with thermal noise $P_{n}=T\delta\omega/2\pi$ with $\delta\omega=10^{-6}m_{a}$,
-$\frac{P(\sigma=10^{-3}\text{eV})}{P_{n}}\simeq 8.6\times 10^{-7}g_{\gamma}^{2}\Big(\frac{20\text{mK}}{T}\Big)\Big(\frac{L}{100\mbox{cm}}\Big)\Big(\frac{R}{6\text{cm}}\Big)^{2}\Big(\frac{B}{15\text{T}}\Big)^{2}\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)$ (4)
+In the observation, the electric $I(\sigma=10^{-3}\text{eV})$ should be compared with thermal noise $I_{n}=\sqrt{2T\delta\omega/\pi R_{c}}$ where the resistance $R_{c}=L/(\sigma\pi R^{2})$ of the cylinder and $\delta\omega=10^{-6}m_{a}$,
+$\frac{I(\sigma=10^{-3}\text{eV})}{I_{n}(\sigma=10^{-3}\text{eV})}\simeq 4.2\times 10^{-4}g_{\gamma}\Big(\frac{20\text{mK}}{T}\Big)^{1/2}\Big(\frac{L}{100\mbox{cm}}\Big)^{1/2}\Big(\frac{R}{6\text{cm}}\Big)\Big(\frac{B}{15\text{T}}\Big)\Big(\frac{10^{-4}\text{eV}}{m_{a}}\Big)^{1/2}\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)^{1/2}$ (4)
 The noise is further reduced by considering a frequency width $\delta\omega=10^{-6}m_{a}$ and an observation time $\delta t_{ob}=100$seconds.
-In this case, the noise decreases by the factor $\sqrt{\delta\omega\delta t_{ob}/2\pi}\simeq 1.6\times 10^{3}$, yielding $P(\sigma=10^{-3}\text{eV})\sqrt{\delta\omega\delta t_{ob}/2\pi}/P_{n}\simeq 1.3\times 10^{-3}$.
-Although the details of power detection are beyond the scope of this paper, we expect the ratio to be sufficiently large for the power ( $P(\sigma=10^{-3},\text{eV})$ ) to become detectable in the near future.
+In this case, the noise decreases by the factor $\sqrt{\delta\omega\delta t_{ob}/2\pi}\simeq 1.6\times 10^{3}$, yielding $I(\sigma=10^{-3}\text{eV})\sqrt{\delta\omega\delta t_{ob}/2\pi}/I_{n}\simeq 6.7\times 10^{-1}$.
+The detection of the current is observed using parallel LC circuit in which the current is amplified with resonance by tuning parameters of the circuit.
+Even when we have small enhancement factor $Q=10$, the signal-noise ratio becomes larger than $1$; $QI(\sigma=10^{-3}\text{eV})\sqrt{\delta\omega\delta t_{ob}/2\pi}/I_{n}\simeq 4.6$.
+Consequently, the signal of the axion can be detected by our method.
 Our method, which employs a cylinder with low electrical conductivity, is also applicable to the detection of the so-called dark photon.
@@
 The large power $P(\sigma=10^{-3}\text{eV})$ arises mainly from large electric current $I$ flowing in the bulk of the cylinder. The electric current $I$ is,
-$I(\sigma=10^{-3}\text{eV})=\frac{\sigma|E_{a}|S}{2\epsilon}\simeq 2.8\times 10^{-14}\text{A}g_{\gamma}\Big(\frac{R}{6\text{cm}}\Big)^{2}\Big(\frac{\sigma}{10^{-3}\text{eV}}\Big)\Big(\frac{B_{0}}{15\text{T}}\Big),$ (27)
+$I(\sigma=10^{-3}\text{eV})=\frac{\sigma|E_{a}|S}{2\epsilon}\simeq 2.8\times 10^{-14}\text{A}g_{\gamma}\Big(\frac{R}{6\text{cm}}\Big)^{2}\Big(\frac{\sigma}{10^{-3}\text{eV}}\Big)\Big(\frac{B_{0}}{15\text{T}}\Big)\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)^{1/2},$ (27)
 where the electric field $E=Re(E^{\prime})=|E_{a}|/2\epsilon=|E_{a}|/20$ is the one in the first term with $y=10$ in eq(16).
@@
 $\Big(\frac{B_{0}}{15\mbox{T}}\Big)^{2}\Big(\frac{m_{a}}{10^{-3}\mbox{eV}}\Big)^{1/2}\Big(\frac{10}{\epsilon}\Big)\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)\Big(\frac{\delta t_{ob}}{100\,\text{s}}\Big)^{1/2}.$ (31)
-To achieve a signal-to-noise ratio greater than 1, the power $P$ must be amplified by approximately three orders of magnitude.
-The method of amplification and its technical implementation are beyond the scope of this paper.
-However, once such amplification is realized, the dark matter axion can be detected using our proposed approach.
+These powers can be detected using a parallel LC circuit, in which electric current in coil is measured.
+By tuning the parameters of the LC circuit, the electric current at resonance is enhanced by a factor $Q=R_{c}/m_{a}L_{c}$ where $L_{c}$ denotes inductance of the inductor and resistance $R_{c}=R$ of the cylinder.
+The cylinder forms the resistor of the circuit.
+The resonance condition is given by $m_{a}=1/\sqrt{C_{c}L_{c}}$ where $C_{c}$ is capacitance of the capacitor.
+In order to observe the electric current $I(\sigma=10^{-3}\text{eV})\simeq 2.8\times 10^{-14}$A, we compare it with the thermal noise $I_{n}$,
+$I_{n}(\sigma=10^{-3}\text{eV})=\sqrt{\frac{2T\delta\omega}{\pi R_{c}}}\simeq 6.6\times 10^{-11}\text{A}\Big(\frac{T}{20\text{mK}}\Big)^{1/2}\Big(\frac{100\text{cm}}{L}\Big)^{1/2}\Big(\frac{R}{6\text{cm}}\Big)\Big(\frac{\sigma}{10^{-3}\text{eV}}\Big)^{1/2}\Big(\frac{m_{a}}{10^{-4}\text{eV}}\Big)^{1/2}$ (32)
+with the resistance $R_{c}=L/(\sigma\pi R^{2})\simeq 1.44(L/100\text{cm})(6\text{cm}/R^{2})\simeq 540\Omega$. Then, a signal-to-noise ratio at resonance is given by
+$\frac{QI(\sigma=10^{-3}\text{eV})}{I_{n}({\sigma=10^{-3}\text{eV})}}$ $\sqrt{\frac{\delta\omega\delta t_{ob}}{2\pi}}\simeq$ (33)
+$0.66Q$ $g_{\gamma}\Big(\frac{20\text{mK}}{T}\Big)^{1/2}\Big(\frac{L}{100\text{cm}}\Big)^{1/2}\Big(\frac{R}{6\text{cm}}\Big)\Big(\frac{B_{0}}{15\mbox{T}}\Big)\Big(\frac{\sigma}{10^{-3}\text{eV}}\Big)^{1/2}\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)^{1/2}\Big(\frac{\delta t_{ob}}{100\,\text{s}}\Big)^{1/2}$
+Consequently, to achieve the signal-to-noise ratio greater than 1, it is sufficient to have $Q\geq 2$.
+Of course, it is favorable to have much large $Q$ value for the axion detection.
+Once such a quality factor $Q$ is realized, the dark matter axion can be detected using our proposed approach.
 IV conclusion
 Summarized, we have proposed a method for axion detection with a cylindrical sample choosing appropriate small conductivity $(\sigma=10^{-3}\text{-}10^{-2}\text{eV})$ sensitive to microwave signals in the frequency range $m_{a}/2\pi=24\text{-}240\text{GHz}$.
-With such conductivity, the axion induces large electric power generated in the bulk of the cylinder and causes a high signal-to-noise ratio.
+With such conductivity, the axion induces large electric current generated in the bulk of the cylinder.
+It causes a high signal-to-noise ratio of the current in parallel LC circuit even with small $Q\geq 2$.
 Therefore, the method demonstrates the feasibility of detecting dark matter axions under realistic experimental conditions.

### v5

@@
 $\sigma=10^{-3}\text{eV}$, the electric current flows inside the bulk of the cylinder.
-Within the QCD axion model, the power $P$ is estimated as $P\simeq 5.7\times 10^{-27}\mbox{W}g_{\gamma}^{2}\Big(\frac{L}{100\mbox{cm}}\Big)\Big(\frac{R}{6\mbox{cm}}\Big)^{2}\Big(\frac{B_{0}}{15\mbox{T}}\Big)^{2}\Big(\frac{m_{a}}{10^{-4}\mbox{eV}}\Big)\Big(\frac{10}{\epsilon}\Big)\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)$ with radius $R$, length $L$, permittivity $\epsilon=10$ of the cylinder and axion energy density $\rho_{a}$, where $g_{\gamma}$ is model dependent parameter; $g_{\gamma}(\text{KSVZ})=-0.96$ and $g_{\gamma}(\text{DFSZ})=0.37$.
-The signal of the axion can be detected by observing electric current in parallel LC circuit with quality factor $Q\geq 2$.
-The detection of dark matter axions in our method may be feasible in the mass range $m_{a}=10^{-4}\text{-}10^{-3}\text{eV}$.
+Within the QCD axion model, the current $I$ is estimated as $I(\sigma=10^{-3}\text{eV})\simeq 2.8\times 10^{-14}\text{A}g_{\gamma}\big(R/6\text{cm}\big)^{2}\big(\sigma/10^{-3}\text{eV}\big)\big(B_{0}/15\text{T}\big)\big(10/\epsilon\big)\big(\rho_{a}/0.3\rm GeVcm^{-3}\big)^{1/2}$ for $m_{a}=10^{-4}$eV, with radius $R$, permittivity $\epsilon=10$ of the cylinder and axion energy density $\rho_{a}$, where $g_{\gamma}$ is model dependent parameter; $g_{\gamma}(\text{KSVZ})=-0.96$ and $g_{\gamma}(\text{DFSZ})=0.37$.
+Because the current is proportional to $R^{2}$, using large sample with R=80R=80cm, we have large signal-noise ratio ( >1>1 ) even in temperature T=4T=4K, $I(\sigma=10^{-3}\text{eV})/I_{n}({\sigma=10^{-3}\text{eV})}\times\sqrt{\delta\omega\delta t_{ob}/2\pi}\simeq 1.1g_{\gamma}(4\text{K}/T)^{1/2}(L/100\text{cm})^{1/2}(R/80\text{cm})(B_{0}/7\mbox{T})(\rho_{a}/0.3\rm GeVcm^{-3})^{1/2}(\delta t_{ob}/10^{3}\,\text{s})^{1/2}$ for $m_{a}=10^{-4}\text{eV}$ with $\epsilon=10$ and $\sigma=\epsilon m_{a}$, where thermal noise is $I_{n}=\sqrt{2T\delta\omega/\pi R_{c}}$ with $\delta\omega=10^{-6}m_{a}$ and resistance $R_{c}=L/(\sigma\pi R^{2})$ of the cylinder with length $L$.
+Although a superconducting solenoid sufficiently large to accommodate such a sample is required, the detection of dark matter axions in our proposal may be feasible in the mass range $m_{a}=10^{-4}\text{-}10^{-3}\text{eV}$.
 I introduction
@@
 The induced electromagnetic radiation is expected to be detected with resonant cavitiesadmx ; haystac , superconducting qubitsmoroi , Quantum Hall effect iwazaki ; aiwazaki , e.t.c.
-In this letter we propose a new method for axion detection using a cylindrical sample with small electrical conductivity, $\sigma=10^{-3}\text{-}10^{-2}\text{eV}$ at low temperature 2020mK.
+In this letter we propose a new method for axion detection using a cylindrical sample with small electrical conductivity, $\sigma=10^{-3}\text{-}10^{-2}\text{eV}$ at low temperature $\sim 1$K.
 Our target is the mass range $m_{a}=10^{-4}\mbox{-}10^{-3}$eV, which is difficult to detect with resonant cavity.
@@
 As a result, the electric current is not large enough to be detectable.
-On the other hand, the two suppression factors do not arise in materials with small conductivity such as semiconductor, Especially, for the cylinder sample with small conductivity $\sigma=10^{-3}$eV, the induced current mainly flows the bulk, and the resulting electric current $I$ and power $P$ become large.
-It would be detectable at low temperatures ( $T\sim 10\text{mK}$ ).
+On the other hand, the two suppression factors do not arise in materials with small conductivity such as semiconductor.
+One is that we have no suppression factor, $\sqrt{m_{a}/\sigma}$ for the cylinder sample with small conductivity $\sigma=10^{-3}$eV.
+The other one is that the induced current mainly flows the bulk.
+The presence of the bulk current is a distinctive feature of electromagnetic fields coupled with axion.
+The resulting electric current $I$ and power $P$ become large with large sample.
+It would be detectable at low temperatures ( $T\sim 1\text{K}$ ).
 We show that when the axion mass $m_{a}=10^{-4}$eV,
-$I(\sigma=10^{-3}\text{eV})$ $\simeq$ $2.8\times 10^{-14}\text{A}g_{\gamma}\Big(\frac{R}{6\text{cm}}\Big)^{2}\Big(\frac{\sigma}{10^{-3}\text{eV}}\Big)\Big(\frac{B_{0}}{15\text{T}}\Big)\left(\frac{10}{\epsilon}\right)\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)^{1/2}$ (1)
-$P(\sigma=10^{-3}\text{eV})$ $\simeq$ $5.7\times 10^{-27}\mbox{W}g_{\gamma}^{2}\Big(\frac{L}{100\mbox{cm}}\Big)\Big(\frac{R}{6\mbox{cm}}\Big)^{2}\Big(\frac{B_{0}}{15\mbox{T}}\Big)^{2}\Big(\frac{m_{a}}{10^{-4}\mbox{eV}}\Big)\Big(\frac{10}{\epsilon}\Big)\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)$ (2)
+$I(\sigma=10^{-3}\text{eV})$ $\simeq$ $2.9\times 10^{-14}\text{A}g_{\gamma}\Big(\frac{R}{6\text{cm}}\Big)^{2}\Big(\frac{B_{0}}{15\text{T}}\Big)\Big(\frac{2\big(\frac{\epsilon}{10}\big)\big(\frac{\sigma}{10^{-3}\text{eV}}\big)\big(\frac{m_{a}}{10^{-4}\text{eV}}\big)^{2}}{\big(\frac{\epsilon}{10}\big)^{2}\big(\frac{m_{a}}{10^{-4}\text{eV}}\big)^{2}+\big(\frac{\sigma}{10^{-3}\text{eV}}\big)^{2}}\Big)\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)^{1/2}$ (1)
+$P(\sigma=10^{-3}\text{eV})$ $\simeq$ $5.3\times 10^{-27}\mbox{W}g_{\gamma}^{2}\Big(\frac{L}{100\mbox{cm}}\Big)\Big(\frac{R}{6\mbox{cm}}\Big)^{2}\Big(\frac{B_{0}}{15\mbox{T}}\Big)^{2}\Big(\frac{2\big(\frac{\sigma}{10^{-3}\text{eV}}\big)\big(\frac{m_{a}}{10^{-4}\text{eV}}\big)^{2}}{\big(\frac{\epsilon}{10}\big)^{2}\big(\frac{m_{a}}{10^{-4}\text{eV}}\big)^{2}+\big(\frac{\sigma}{10^{-3}\text{eV}}\big)^{2}}\Big)\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)$ (2)
 with the choice $\sigma=10m_{a}=10^{-3}$eV, where $R(L)$ denotes the radius ( length ) of the cylinder with permittivity $\epsilon=10$.
@@
 Indeed, when the conductivity is large, for instance, $\sigma=10^{4}$eV, the power is given by
-$P(\sigma=10^{4}\text{eV})\simeq 5.4\times 10^{-31}\mbox{W}g_{\gamma}^{2}\Big(\frac{L}{100\mbox{cm}}\Big)\Big(\frac{R}{6\mbox{cm}}\Big)\sqrt{\frac{m_{a}}{10^{-4}\mbox{eV}}}\sqrt{\frac{10^{4}\mbox{eV}}{\sigma}}\Big(\frac{B}{15\mbox{T}}\Big)^{2}\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)$ (3)
+$P(\sigma=10^{4}\text{eV})\simeq 5.4\times 10^{-31}\mbox{W}g_{\gamma}^{2}\Big(\frac{L}{100\mbox{cm}}\Big)\Big(\frac{R}{6\mbox{cm}}\Big)\sqrt{\frac{m_{a}}{10^{-4}\mbox{eV}}}\sqrt{\frac{10^{4}\mbox{eV}}{\sigma}}\Big(\frac{B_{0}}{15\mbox{T}}\Big)^{2}\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)$ (3)
 Obviously, the power $P(\sigma=10^{4}\text{eV})$ proportional to $R$ is fourth order of magnitude smaller than $P(\sigma=10^{-3}\text{eV})$.
-In the observation, the electric $I(\sigma=10^{-3}\text{eV})$ should be compared with thermal noise $I_{n}=\sqrt{2T\delta\omega/\pi R_{c}}$ where the resistance $R_{c}=L/(\sigma\pi R^{2})$ of the cylinder and $\delta\omega=10^{-6}m_{a}$,
-$\frac{I(\sigma=10^{-3}\text{eV})}{I_{n}(\sigma=10^{-3}\text{eV})}\simeq 4.2\times 10^{-4}g_{\gamma}\Big(\frac{20\text{mK}}{T}\Big)^{1/2}\Big(\frac{L}{100\mbox{cm}}\Big)^{1/2}\Big(\frac{R}{6\text{cm}}\Big)\Big(\frac{B}{15\text{T}}\Big)\Big(\frac{10^{-4}\text{eV}}{m_{a}}\Big)^{1/2}\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)^{1/2}$ (4)
-The noise is further reduced by considering a frequency width $\delta\omega=10^{-6}m_{a}$ and an observation time $\delta t_{ob}=100$seconds.
-In this case, the noise decreases by the factor $\sqrt{\delta\omega\delta t_{ob}/2\pi}\simeq 1.6\times 10^{3}$, yielding $I(\sigma=10^{-3}\text{eV})\sqrt{\delta\omega\delta t_{ob}/2\pi}/I_{n}\simeq 6.7\times 10^{-1}$.
-The detection of the current is observed using parallel LC circuit in which the current is amplified with resonance by tuning parameters of the circuit.
-Even when we have small enhancement factor $Q=10$, the signal-noise ratio becomes larger than $1$; $QI(\sigma=10^{-3}\text{eV})\sqrt{\delta\omega\delta t_{ob}/2\pi}/I_{n}\simeq 4.6$.
-Consequently, the signal of the axion can be detected by our method.
-Our method, which employs a cylinder with low electrical conductivity, is also applicable to the detection of the so-called dark photon.
-The electromagnetic coupling of the dark photon produces an effect similar to that of the axion.
+In the observation, the electric current $I(\sigma=10^{-3}\text{eV})$ should be compared with thermal noise $I_{n}=\sqrt{2T\delta\omega/\pi R_{c}}$ where the resistance $R_{c}=L/(\sigma\pi R^{2})$ of the cylinder and $\delta\omega=10^{-6}m_{a}$,
+$\frac{I\big(\sigma=(10^{-3}\text{eV}/\sqrt{3}\big))}{I_{n}(\sigma=\big(10^{-3}\text{eV}/\sqrt{3})\big)}$ $\simeq$ $5.0\times 10^{-4}g_{\gamma}\Big(\frac{20\text{mK}}{T}\Big)^{1/2}\Big(\frac{L}{100\mbox{cm}}\Big)^{1/2}\Big(\frac{R}{6\text{cm}}\Big)$ (4)
+$\times$ $\Big(\frac{B_{0}}{15\text{T}}\Big)\Big(\frac{2\big(\frac{\epsilon}{10}\big)\big(\frac{\sqrt{3}\sigma}{10^{-3}\text{eV}}\big)^{1/2}\big(\frac{m_{a}}{10^{-4}\text{eV}}\big)^{3/2}}{\big(\frac{\epsilon}{10}\big)^{2}\big(\frac{m_{a}}{10^{-4}\text{eV}}\big)^{2}+\big(\frac{\sqrt{3}\sigma}{10^{-3}\text{eV}}\big)^{2}}\Big)\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)^{1/2},$
+with $m_{a}=10^{-4}$eV, where we have taken $\sigma=\epsilon m_{a}/\sqrt{3}=(10^{-3}\text{eV}/\sqrt{3})$, which maximizes the ratio $I/I_{n}$.
+The noise is further reduced by considering a frequency width $\delta\omega=10^{-6}m_{a}$ and an observation time $\delta t_{ob}=10^{3}$seconds.
+In this case, the noise decreases by the factor $\sqrt{\delta\omega\delta t_{ob}/2\pi}\simeq 4.9\times 10^{3}$, yielding
+$\frac{I\big(\sigma=(10^{-3}\text{eV}/\sqrt{3})\big)\sqrt{\frac{\delta\omega\delta t_{ob}}{2\pi}}}{I_{n}\big(\sigma=(10^{-3}\text{eV}/\sqrt{3})\big)}\simeq 2.5.$ (5)
+Achieving an electrical conductivity of order $10^{-3}$eV in semiconductors at 2020mK is challenging, but becomes feasible at 44K.
+Although thermal noise is higher at this temperature, the signal can be correspondingly enhanced when we take large radius $R$.
+Since the induced current scales with the square of the sample radius, the use of larger samples, i.e.
+R=50R=50cm, significantly increases the signal.
+Consequently, a large signal-to-noise ratio can be achieved even at 44K.
+$\frac{I\big(\sigma=\big(10^{-3}\text{eV}/\sqrt{3})\big)}{I_{n}\big({\sigma=(10^{-3}\text{eV}/\sqrt{3})\big)}}$ $\times$ $\sqrt{\frac{\delta\omega\delta t_{ob}}{2\pi}}\simeq 0.98g_{\gamma}\Big(\frac{4\text{K}}{T}\Big)^{1/2}\Big(\frac{L}{100\text{cm}}\Big)^{1/2}\Big(\frac{R}{50\text{cm}}\Big)$ (6)
+$\times$ $\Big(\frac{B_{0}}{10\mbox{T}}\Big)\Big(\frac{2\big(\frac{\epsilon}{10}\big)\big(\frac{\sqrt{3}\sigma}{10^{-3}\text{eV}}\big)^{1/2}\big(\frac{m_{a}}{10^{-4}\text{eV}}\big)^{2}}{\big(\frac{\epsilon}{10}\big)^{2}\big(\frac{m_{a}}{10^{-4}\text{eV}}\big)^{2}+\big(\frac{\sqrt{3}\sigma}{10^{-3}\text{eV}}\big)^{2}}\Big)\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)^{1/2}\Big(\frac{\delta t_{ob}}{1000\,\text{s}}\Big)^{1/2}.$
+Our method, which employs a cylinder with low electrical conductivity, is also applicable to the detection of axion like particle ( ALP ) or so-called dark photon.
+The electromagnetic coupling of theses particles produces a signal similar to that of the axion.
+In the next section, we briefly explain the property of QDC axion coupled with electromagnetic fields.
+Using the coupling, we solve electromagnetic fields generated inside the cylinder under strong magnetic field and calculate electric power $P$ and current $I$ in section (III).
+In the section (IV), we discuss realistic choices of materials forming the large cylinder R=80R=80cm and temperature 44K cooling the cylinder.
+We present a figure in which detectable regions of ALPs are shown in $g_{a\gamma\gamma}-m_{a}$ plane.
+We also show that ALPs with extremely small mass, e.g.
+$10^{-14}$eV can be detectable with the choice of electric conductivity being of the order of the mass.
 II axion
@@
 In the limit, the main contribution comes from the second term in $U$.
-We note that $|H_{1}^{(1)}(x)/H_{0}^{(1)}(x)|^{2}\simeq 1$ for $x>5$.
+We note that $|H_{1}^{(1)}(x)/H_{0}^{(1)}(x)|^{2}\simeq 1$ for $x>5$ and that $b(y)\to\sqrt{\sigma/m_{a}}\exp(i\pi/4)$ in the limit $y=\sigma/m_{a}\gg 1$.
 The skin depth for general $\sigma$ represented by the second term of $U(x,y,z)$ in eq(22) is given by $\big(m_{a}\mbox{Im}(b(y))\big)^{-1}=\Big(m_{a}(\epsilon^{2}+(\sigma/m_{a})^{2})^{1/4}\sin(\theta/2)\Big)^{-1}>\delta=\sqrt{2/m_{a}\sigma}$.
@@
 But when we put $\sigma=\epsilon m_{a}$ in the formula, the dependence of $P$ on $m_{a}$ becomes simple; $P\propto m_{a}$.
-In the actual search of the axion mass in the range $10^{-4}\text{-}10^{-3}$eV, it is sufficient for the axion detection to take the value $\sigma\sim 5\times 10^{-3}$eV with $\epsilon\sim 10$.
+Otherwise, the dependence of $P$ on $m_{a}$ or $\sigma$ is not simply scaling; $P\propto\sigma m_{a}^{2}/(\epsilon^{2}m_{a}^{2}+\sigma^{2})$.
+In the actual search of the axion mass in the range $10^{-4}\text{-}10^{-3}$eV, it is sufficient for the axion detection to take the value $\sigma\sim 1.0\times 10^{-3}$eV with $\epsilon\sim 10$.
 Although it does not lead to the maximal power $P$, the predicted signal-noise ratios are large enough for the detection.
-Hereafter we take $\sigma=\epsilon m_{a}$ and $\epsilon=10$ for simplicity.
+Hereafter, we take $\sigma=\epsilon m_{a}$ and $\epsilon=10$ to estimate the power $P$.
+Later we find that $\sigma=\epsilon m_{a}/\sqrt{3}$ maximizes the ratio of electric current $I$ induced by axion to thermal noise $I_{n}$.
+Please refer Fig.2 in which the dependence of $P(\sigma,m_{a})$ and the current $I(\sigma,m_{a})$ on $\sigma$ and $m_{a}$ is shown.
 Numerically, noting $|E_{a}|=|g_{a\gamma\gamma}a(t)B_{0}|\simeq 2.4\times 10^{-36}\text{GeV}^{2}g_{\gamma}(B_{0}/15\text{T})$, we find
-$P(\sigma=10^{-3}\text{eV})\simeq 5.7\times 10^{-27}\mbox{W}g_{\gamma}^{2}\Big(\frac{L}{100\mbox{cm}}\Big)\Big(\frac{R}{6\mbox{cm}}\Big)^{2}\Big(\frac{B_{0}}{15\mbox{T}}\Big)^{2}\Big(\frac{m_{a}}{10^{-4}\mbox{eV}}\Big)\Big(\frac{10}{\epsilon}\Big)\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)$ (24)
-with $\epsilon=10$, $y=10$ ( $\sigma=10m_{a}$ ) and $x\simeq 30$ ( R=6R=6cm ).
+$P(\sigma=10^{-3}\text{eV})\simeq 5.3\times 10^{-27}\mbox{W}g_{\gamma}^{2}\Big(\frac{L}{100\mbox{cm}}\Big)\Big(\frac{R}{6\mbox{cm}}\Big)^{2}\Big(\frac{B_{0}}{15\mbox{T}}\Big)^{2}\Big(\frac{2\big(\frac{\sigma}{10^{-3}\text{eV}}\big)\big(\frac{m_{a}}{10^{-4}\text{eV}}\big)^{2}}{\big(\frac{\epsilon}{10}\big)^{2}\big(\frac{m_{a}}{10^{-4}\text{eV}}\big)^{2}+\big(\frac{\sigma}{10^{-3}\text{eV}}\big)^{2}}\Big)\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)$ (26)
+with $y=10$ ( $\sigma=10m_{a}$ ) and $x\simeq 30$ ( $x=m_{a}R=10^{-3}\text{eV}\times 6\text{cm}\simeq 30$ ).
 Obviously, $P(\sigma=10^{-3}\text{eV})$ in eq.(26) is proportional to $R^{2}=x^{2}/m_{a}^{2}$, that is, the cross section $\pi R^{2}$ of the cylinder.
@@
 The large power $P(\sigma=10^{-3}\text{eV})$ arises mainly from large electric current $I$ flowing in the bulk of the cylinder. The electric current $I$ is,
-$I(\sigma=10^{-3}\text{eV})=\frac{\sigma|E_{a}|S}{2\epsilon}\simeq 2.8\times 10^{-14}\text{A}g_{\gamma}\Big(\frac{R}{6\text{cm}}\Big)^{2}\Big(\frac{\sigma}{10^{-3}\text{eV}}\Big)\Big(\frac{B_{0}}{15\text{T}}\Big)\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)^{1/2},$ (27)
-where the electric field $E=Re(E^{\prime})=|E_{a}|/2\epsilon=|E_{a}|/20$ is the one in the first term with $y=10$ in eq(16).
+$I(\sigma=10^{-3}\text{eV})=\frac{\sigma\epsilon|E_{a}|S}{(\epsilon^{2}+y^{2})}\simeq 2.9\times 10^{-14}\text{A}g_{\gamma}\Big(\frac{R}{6\text{cm}}\Big)^{2}\Big(\frac{2\big(\frac{\epsilon}{10}\big)\big(\frac{\sigma}{10^{-3}\text{eV}}\big)\big(\frac{m_{a}}{10^{-4}\text{eV}}\big)^{2}}{\big(\frac{\epsilon}{10}\big)^{2}\big(\frac{m_{a}}{10^{-4}\text{eV}}\big)^{2}+\big(\frac{\sigma}{10^{-3}\text{eV}}\big)^{2}}\Big)\Big(\frac{B_{0}}{15\text{T}}\Big)\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)^{1/2},$ (29)
+where the electric field $E=Re(E^{\prime})=\epsilon|E_{a}|/(\epsilon^{2}+y^{2})=|E_{a}|/20$ is the one in the first term with $y=10$ in eq(18).
 The point causing the large current $I(\sigma=10^{-3}\text{eV})$ is that electric field $E$ inside of the cylinder with small conductivity, e.g.
 $\sigma=10^{-3}$eV has no suppression factor $\sqrt{m_{a}/\sigma}\simeq 10^{-4}$, which is present when the cylinder with large conductivity $\sigma=10^{4}$eV.
-Additionally, the current flows in the whole of the cylinder, not just in the surface.
+Additionally, the current flows in the bulk of the cylinder, not just in the surface.
 The facts make the large electric current $I(\sigma=10^{-3}\text{eV})\sim 10^{-14}$A.
+Figure 2 shows the dependence of the power $P(\sigma,m_{a})$ and current $I(\sigma,m_{a})$ on the axion mass $m_{a}$ and electrical conductivity $\sigma$ when $\epsilon=10$.
+The vertical axis is given in arbitrary units.
+It is remarkable feature that as the axion mass $m_{a}$ increases, increasing the electrical conductivity $\sigma$ correspondingly leads to larger values of both $P(\sigma,m_{a})$ and $I(\sigma,m_{a})$.
+Figure 2: Dependence of power $P(\sigma,m_{a})$ and current $I(\sigma,m_{a})$ on $\sigma$ and $m_{a}$ when $\epsilon=10$.
 In order to observe the power $P(\sigma=10^{-3}\text{eV})$, we need to take into account thermal noise ( Johnson-Nyquist noise ).
@@
 $\frac{P(\sigma=10^{-3}\text{eV})}{P_{n}}\sqrt{\frac{\delta\omega\delta t_{ob}}{2\pi}}\simeq 1.3\times 10^{-3}g_{\gamma}^{2}\Big(\frac{20\text{mK}}{T}\Big)\Big(\frac{L}{100\mbox{cm}}\Big)\Big(\frac{R}{6\mbox{cm}}\Big)^{2}\times$
-$\Big(\frac{B_{0}}{15\mbox{T}}\Big)^{2}\Big(\frac{m_{a}}{10^{-4}\mbox{eV}}\Big)^{1/2}\Big(\frac{10}{\epsilon}\Big)\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)\Big(\frac{\delta t_{ob}}{100\,\text{s}}\Big)^{1/2}.$ (30)
+$\Big(\frac{B_{0}}{15\mbox{T}}\Big)^{2}\Big(\frac{2\big(\frac{\sigma}{10^{-3}\text{eV}}\big)\big(\frac{m_{a}}{10^{-4}\text{eV}}\big)^{3/2}}{\big(\frac{\epsilon}{10}\big)^{2}\big(\frac{m_{a}}{10^{-4}\text{eV}}\big)^{2}+\big(\frac{\sigma}{10^{-3}\text{eV}}\big)^{2}}\Big)\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)\Big(\frac{\delta t_{ob}}{100\,\text{s}}\Big)^{1/2}.$ (32)
 It is the ratio when $m_{a}=10^{-4}$eV and $\sigma=10^{-3}$eV. Further, when $m_{a}=10^{-3}$eV, taking $\sigma=10m_{a}=10^{-2}$eV, we have
 $\frac{P(\sigma=10^{-2}\text{eV})}{P_{n}}\sqrt{\frac{\delta\omega\delta t_{ob}}{2\pi}}\simeq 4.1\times 10^{-3}g_{\gamma}^{2}\Big(\frac{20\text{mK}}{T}\Big)\Big(\frac{L}{100\mbox{cm}}\Big)\Big(\frac{R}{6\mbox{cm}}\Big)^{2}\times$
-$\Big(\frac{B_{0}}{15\mbox{T}}\Big)^{2}\Big(\frac{m_{a}}{10^{-3}\mbox{eV}}\Big)^{1/2}\Big(\frac{10}{\epsilon}\Big)\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)\Big(\frac{\delta t_{ob}}{100\,\text{s}}\Big)^{1/2}.$ (31)
-These powers can be detected using a parallel LC circuit, in which electric current in coil is measured.
-By tuning the parameters of the LC circuit, the electric current at resonance is enhanced by a factor $Q=R_{c}/m_{a}L_{c}$ where $L_{c}$ denotes inductance of the inductor and resistance $R_{c}=R$ of the cylinder.
-The cylinder forms the resistor of the circuit.
-The resonance condition is given by $m_{a}=1/\sqrt{C_{c}L_{c}}$ where $C_{c}$ is capacitance of the capacitor.
-In order to observe the electric current $I(\sigma=10^{-3}\text{eV})\simeq 2.8\times 10^{-14}$A, we compare it with the thermal noise $I_{n}$,
+$\Big(\frac{B_{0}}{15\mbox{T}}\Big)^{2}\Big(\frac{2\big(\frac{\sigma}{10^{-2}\text{eV}}\big)\big(\frac{m_{a}}{10^{-3}\text{eV}}\big)^{3/2}}{\big(\frac{\epsilon}{10}\big)^{2}\big(\frac{m_{a}}{10^{-3}\text{eV}}\big)^{2}+\big(\frac{\sigma}{10^{-2}\text{eV}}\big)^{2}}\Big)\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)\Big(\frac{\delta t_{ob}}{100\,\text{s}}\Big)^{1/2}.$ (33)
+SN ratio of the power is less than $10^{-2}$, but we can show that SN ratio of electric current $I$ is bigger than $1$.
+We should make a comment that we have taken $y\equiv\sigma/m_{a}=\epsilon$ in order to maximize the SN ratio $P/P_{n}$.
+On the other hand, to maximize SN ratio $I/I_{n}$ of the electric current to thermal noise $I_{n}$, we need to take $y=\epsilon/\sqrt{3}$, because $I/I_{n}\propto yy^{-1/2}/(\epsilon^{2}+y^{2})$: $I_{n}=\sqrt{2T\delta\omega/\pi R_{c}}$ with the resistance of the cylinder, $R_{c}=L/(\sigma\pi R^{2})$.
+However, we take $y=\epsilon$ for simplicity in the subsequent calculations.
+It simply causes slightly smaller value of the ratio $I/I_{n}$; $y^{1/2}/(\epsilon^{2}+y^{2})\simeq 1.58\times 10^{-3}$ with $y=\epsilon=10$, while $y^{1/2}/(\epsilon^{2}+y^{2})\simeq 1.8\times 10^{-3}$ with $y=\epsilon/\sqrt{3}$.
+Because thermal noise $I_{n}$ is
 $I_{n}(\sigma=10^{-3}\text{eV})=\sqrt{\frac{2T\delta\omega}{\pi R_{c}}}\simeq 6.9\times 10^{-11}\text{A}\Big(\frac{T}{20\text{mK}}\Big)^{1/2}\Big(\frac{100\text{cm}}{L}\Big)^{1/2}\Big(\frac{R}{6\text{cm}}\Big)\Big(\frac{\sigma}{10^{-3}\text{eV}}\Big)^{1/2}\Big(\frac{m_{a}}{10^{-4}\text{eV}}\Big)^{1/2}$ (34)
-with the resistance $R_{c}=L/(\sigma\pi R^{2})\simeq 1.44(L/100\text{cm})(6\text{cm}/R^{2})\simeq 540\Omega$. Then, a signal-to-noise ratio at resonance is given by
-$\frac{QI(\sigma=10^{-3}\text{eV})}{I_{n}({\sigma=10^{-3}\text{eV})}}$ $\sqrt{\frac{\delta\omega\delta t_{ob}}{2\pi}}\simeq$ (33)
-$0.66Q$ $g_{\gamma}\Big(\frac{20\text{mK}}{T}\Big)^{1/2}\Big(\frac{L}{100\text{cm}}\Big)^{1/2}\Big(\frac{R}{6\text{cm}}\Big)\Big(\frac{B_{0}}{15\mbox{T}}\Big)\Big(\frac{\sigma}{10^{-3}\text{eV}}\Big)^{1/2}\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)^{1/2}\Big(\frac{\delta t_{ob}}{100\,\text{s}}\Big)^{1/2}$
-Consequently, to achieve the signal-to-noise ratio greater than 1, it is sufficient to have $Q\geq 2$.
-Of course, it is favorable to have much large $Q$ value for the axion detection.
-Once such a quality factor $Q$ is realized, the dark matter axion can be detected using our proposed approach.
-IV conclusion
-Summarized, we have proposed a method for axion detection with a cylindrical sample choosing appropriate small conductivity $(\sigma=10^{-3}\text{-}10^{-2}\text{eV})$ sensitive to microwave signals in the frequency range $m_{a}/2\pi=24\text{-}240\text{GHz}$.
-With such conductivity, the axion induces large electric current generated in the bulk of the cylinder.
-It causes a high signal-to-noise ratio of the current in parallel LC circuit even with small $Q\geq 2$.
-Therefore, the method demonstrates the feasibility of detecting dark matter axions under realistic experimental conditions.
+with the resistance $R_{c}\simeq 1.44(L/100\text{cm})(6\text{cm}/R)^{2}(10^{-3}\text{eV}/\sigma)\simeq 540\Omega$, the SN ratio is
+$\frac{I(\sigma=10^{-3}\text{eV})}{I_{n}({\sigma=10^{-3}\text{eV})}}$ $\sqrt{\frac{\delta\omega\delta t_{ob}}{2\pi}}$ $\simeq 2.1g_{\gamma}\Big(\frac{20\text{mK}}{T}\Big)^{1/2}\Big(\frac{L}{100\text{cm}}\Big)^{1/2}\Big(\frac{R}{6\text{cm}}\Big)$ (35)
+$\times\Big(\frac{B_{0}}{15\text{T}}\Big)$ $\Big(\frac{2\big(\frac{\epsilon}{10}\big)\big(\frac{\sigma}{10^{-3}\text{eV}}\big)^{1/2}\big(\frac{m_{a}}{10^{-4}\text{eV}}\big)^{2}}{\big(\frac{\epsilon}{10}\big)^{2}\big(\frac{m_{a}}{10^{-4}\text{eV}}\big)^{2}+\big(\frac{\sigma}{10^{-3}\text{eV}}\big)^{2}}\Big)\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)^{1/2}\Big(\frac{\delta t_{ob}}{1000\,\text{s}}\Big)^{1/2}$
+It is the ratio corresponding to axion mass $m_{a}=10^{-4}$eV, while it is given by
+$\frac{I(\sigma=10^{-2}\text{eV})}{I_{n}({\sigma=10^{-2}\text{eV})}}\times$ $\sqrt{\frac{\delta\omega\delta t_{ob}}{2\pi}}$ $\simeq 6.6g_{\gamma}\Big(\frac{20\text{mK}}{T}\Big)^{1/2}\Big(\frac{L}{100\text{cm}}\Big)^{1/2}\Big(\frac{R}{6\text{cm}}\Big)$ (36)
+$\times$ $\Big(\frac{B_{0}}{15\mbox{T}}\Big)\Big(\frac{2\big(\frac{\epsilon}{10}\big)\big(\frac{\sigma}{10^{-2}\text{eV}}\big)^{1/2}\big(\frac{m_{a}}{10^{-3}\text{eV}}\big)^{2}}{\big(\frac{\epsilon}{10}\big)^{2}\big(\frac{m_{a}}{10^{-3}\text{eV}}\big)^{2}+\big(\frac{\sigma}{10^{-2}\text{eV}}\big)^{2}}\Big)\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)^{1/2}\Big(\frac{\delta t_{ob}}{1000\,\text{s}}\Big)^{1/2}$
+for $m_{a}=10^{-3}$eV.
+Consequently, we find that the axion with mass $m_{a}=10^{-4}\text{-}10^{-3}\text{eV}$ can be detected by observing the electric current induced in the semiconductor cylinder.
+Because their frequencies are large such as $24\text{-}240$GHz, it may be detected with heterodyne detector.
+IV discussion
+Increasing the radius $R$ of the cylindrical sample enhances the induced current $I\propto R^{2}$ and thus improves detectability.
+However, in a dilution refrigerator operating at $T=20$ mK, the allowable sample size is severely limited.
+In contrast, at $T=4$ K, large samples can be readily cooled by direct immersion in liquid He.
+By applying a strong magnetic field to such large samples using a superconducting solenoid, the axion-induced current can be significantly enhanced, facilitating detection.
+Moreover, at $20$ mK, achieving a conductivity of order $10^{-3}$eV is challenging for semiconductors and may require alternative materials such as amorphous systems.
+At $4$ K, however, semiconductors with appropriate doping can readily meet this requirement and serve as suitable samples.
+For example, we dope phosphorus into silicon at a critical concentration of approximately $\sim 3.5\times 10^{18}\text{cm}^{-3}$, beyond which the semiconductor becomes a metal.
+Adjacent localized wave functions of electrons doped overlap, creating a new conduction band.
+It leads to the metal-insulator transition.
+Thus, we adjust a carrier mobility $\mu\sim 10\text{cm}^{2}/\text{Vs}$ and may obtain $\sigma=10^{-3}-10^{-2}\text{eV}$.
+We remember that the conductivity $\sigma$ is approximately given by $\sigma=en\mu$; $n$ is carrier density.
+It is easier to perform such an adjustment at 44K than 2020mK.
+Taking into account a detectable signal-to-noise ratio, the following experimental setup can be considered. That is, large sample scale is L=100L=100cm and R=80R=80cm, and magnetic field B=7B=7T.
+$\frac{I\big(\sigma=(10^{-3}\text{eV}/\sqrt{3})\big)}{I_{n}({\sigma=\big(10^{-3}\text{eV}/\sqrt{3})\big)}}$ $\times$ $\sqrt{\frac{\delta\omega\delta t_{ob}}{2\pi}}\simeq 1.1g_{\gamma}\Big(\frac{4\text{K}}{T}\Big)^{1/2}\Big(\frac{L}{100\text{cm}}\Big)^{1/2}$ (37)
+$\times$ $\Big(\frac{R}{80\text{cm}}\Big)\Big(\frac{B_{0}}{7\mbox{T}}\Big)\Big(\frac{2\big(\frac{\epsilon}{10}\big)\big(\frac{\sqrt{3}\sigma}{10^{-3}\text{eV}}\big)^{1/2}\big(\frac{m_{a}}{10^{-4}\text{eV}}\big)^{2}}{\big(\frac{\epsilon}{10}\big)^{2}\big(\frac{m_{a}}{10^{-4}\text{eV}}\big)^{2}+\big(\frac{\sqrt{3}\sigma}{10^{-3}\text{eV}}\big)^{2}}\Big)\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)^{1/2}\Big(\frac{\delta t_{ob}}{1000\,\text{s}}\Big)^{1/2}$
+for $m_{a}=10^{-4}$eV, where we have taken $\sigma=\epsilon m_{a}/\sqrt{3}$ to maximize the ratio. The electric current is large such as $I\simeq 2.1\times 10^{-12}$A.
+Similarly,
+$\frac{I\big(\sigma=(10^{-2}\text{eV}/\sqrt{3})\big)}{I_{n}\big({\sigma=(10^{-2}\text{eV}/\sqrt{3})\big)}}$ $\times$ $\sqrt{\frac{\delta\omega\delta t_{ob}}{2\pi}}\simeq 3.4g_{\gamma}\Big(\frac{4\text{K}}{T}\Big)^{1/2}\Big(\frac{L}{100\text{cm}}\Big)^{1/2}$ (38)
+$\times$ $\Big(\frac{R}{80\text{cm}}\Big)\Big(\frac{B_{0}}{7\mbox{T}}\Big)\Big(\frac{2\big(\frac{\epsilon}{10}\big)\big(\frac{\sqrt{3}\sigma}{10^{-2}\text{eV}}\big)^{1/2}\big(\frac{m_{a}}{10^{-3}\text{eV}}\big)^{2}}{\big(\frac{\epsilon}{10}\big)^{2}\big(\frac{m_{a}}{10^{-3}\text{eV}}\big)^{2}+\big(\frac{\sqrt{3}\sigma}{10^{-2}\text{eV}}\big)^{2}}\Big)\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)^{1/2}\Big(\frac{\delta t_{ob}}{1000\,\text{s}}\Big)^{1/2}$
+for $m_{a}=10^{-3}$eV. The electric current is $I\simeq 2.1\times 10^{-11}$A.
+Obviously, we need a large superconducting solenoid to incorporate the large sample with L=100L=100cm and R=80R=80cm.
+But, the sample volume in eq(37) and eq(38) is too large to be efficiently cooled and to be incorporated in a superconducting solenoid, one can instead use ten smaller samples, each with a radius of 8 cm and a length of 10 cm, connected in parallel.
+In such a configuration, the use of multiple superconducting solenoids, possibly on the order of ten, may be required instead of a single magnet.
+Each one incorporates a small sample.
+Because the samples are connected in parallel, total electric current is identical to the one generated in a large sample.
+Then, a comparable signal-to-noise ratio can be achieved.
+In this way, if a large superconducting solenoid capable of accommodating a sample with L=100L=100cm and RR=80cm, or alternatively several smaller superconducting solenoids accommodating samples with L=10L=10cm and RR=8 cm connected in parallel, can be realized, then the detection of axions with masses in the range $m_{a}=10^{-4}-10^{-3}$eV is feasible.
+Up to now, we have presented results in QCD axion model. More generally, we show SN ratio of electric current in terms of $g_{a\gamma\gamma}$ and $m_{a}$,
+$\frac{I\big(\sigma=(10^{-3}\text{eV}/\sqrt{3})\big)}{I_{n}({\sigma=\big(10^{-3}\text{eV}/\sqrt{3})\big)}}$ $\times$ $\sqrt{\frac{\delta\omega\delta t_{ob}}{2\pi}}\simeq 2.9\times 10^{-2}\Big(\frac{g_{a\gamma\gamma}}{10^{-15}\text{GeV}^{-1}}\Big)\Big(\frac{4\text{K}}{T}\Big)^{1/2}\Big(\frac{L}{100\text{cm}}\Big)^{1/2}$ (39)
+$\times$ $\Big(\frac{R}{80\text{cm}}\Big)\Big(\frac{B_{0}}{7\mbox{T}}\Big)\Big(\frac{2\big(\frac{\epsilon}{10}\big)\big(\frac{\sqrt{3}\sigma}{10^{-3}\text{eV}}\big)^{1/2}\big(\frac{m_{a}}{10^{-4}\text{eV}}\big)}{\big(\frac{\epsilon}{10}\big)^{2}\big(\frac{m_{a}}{10^{-4}\text{eV}}\big)^{2}+\big(\frac{\sqrt{3}\sigma}{10^{-3}\text{eV}}\big)^{2}}\Big)\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)^{1/2}\Big(\frac{\delta t_{ob}}{1000\,\text{s}}\Big)^{1/2}$
+Below, we have presented the figure where the boundaries between SN ratio $I/I_{n}\times\sqrt{\delta\omega\delta t_{ob}/2\pi}>1$ and the ratio <1<1 are shown in $g_{a\gamma\gamma}-m_{a}$ plane.
+It shows that for larger axion masses, increasing the electrical conductivity is beneficial for the detection.
+Figure 3: Boundaries above which SN ratio $I/I_{n}\times\sqrt{\delta\omega\delta t_{ob}/2\pi}$ is larger than 11 for two types of conductivity $\sigma=10^{-3}\text{eV}/\sqrt{3}$ ( dashed ) and $\sigma=10^{-2}\text{eV}/\sqrt{3}$ ( thick ), when $L=100\text{cm}$, $R=80\text{cm}$ and $B_{0}=7$T.
+The ratio increases more as $g_{a\gamma\gamma}$ is bigger above the boundaries.
+We should mention a distinctive feature of our proposal for axion like particle ( ALP ) detection.
+When the mass of the axion is extremely small $m_{a}\sim 10^{-14}$eV, it is detectable for even small sample with L=10L=10cm and R=1R=1cm by taking small conductivity $\sigma\sim 10^{-13}\text{eV}$: The SN ratio is given by $\sim 10$.
+That is, ALPs with extremely small mass are detectable using sample with extremely small conductivity being of the order of the mass, without the use of large superconducting magnets to incorporate the sample.
+It can be understood by noting that the ratio $I/I_{n}\times\sqrt{\delta\omega\delta t_{ob}/2\pi}$ scales such as $\lambda^{-1/2}$ for $m_{a}\to m_{a}\lambda$ and $\sigma\to\sigma\lambda$.
+Thus, smaller $\lambda$ leads to larger ratio.
+V conclusion
+In a strong magnetic field, axions generate an electric current in a cylindrical semiconductor sample.
+This current consists of two components: one localized near the surface due to conventional electromagnetic induction, and another that flows throughout the bulk of the semiconductor.
+The presence of this bulk current is a distinctive feature of axion-induced electromagnetism.
+In semiconductors, whose electrical conductivity is lower than that of metals but significantly higher than that of insulators, the bulk current becomes the dominant contribution.
+In this paper, we have proposed a method for axion detection with a cylindrical sample choosing appropriate small conductivity $(\sigma=10^{-3}\text{-}10^{-2}\text{eV})$, which is sensitive to microwave signals in the frequency range $m_{a}/2\pi=24\text{-}240\text{GHz}$.
+With such conductivity, the axion induces large electric current $I$ in the bulk of the cylinder.
+It causes a high signal-to-noise ratio of the current $I\sqrt{\delta\omega\delta t_{ob}/2\pi}/I_{n}>1$ with the observation time $\delta t_{ob}=10^{3}$second.
+Therefore, the method demonstrates the feasibility of detecting dark matter axions ( or ALPs ) under realistic experimental conditions.
 It is also effective for the search of dark photon.